### v1

# Multi-Modal Attention Networks with Uncertainty Quantification for Automated Concrete Bridge Deck Delamination Detection

Abstract
Deteriorating civil infrastructure requires automated inspection techniques that overcome limitations of traditional visual assessment. While Ground Penetrating Radar and Infrared Thermography enable subsurface defect detection, single-modal approaches face complementary constraints—radar struggles with moisture interference and shallow defects, while thermography exhibits weather dependency and limited penetration depth. This paper presents a multi-modal attention network that fuses radar temporal patterns with thermal spatial signatures for automated bridge deck delamination detection. Our architecture introduces temporal attention for radar signal processing, spatial attention for thermal feature extraction, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns invisible to individual sensors. We incorporate uncertainty quantification through Monte Carlo dropout and learned variance estimation, decomposing predictive uncertainty into epistemic and aleatoric components essential for safety-critical decisions. Experiments on five bridge datasets reveal that on balanced to moderately imbalanced data, our approach substantially outperforms baselines in accuracy and area under curve—representing meaningful improvements over single-modal approaches and concatenation-based fusion. Ablation studies demonstrate that cross-modal attention provides critical gains beyond within-modality attention alone, while multi-head mechanisms achieve improved calibration over single-head alternatives. Uncertainty quantification reduces calibration error, enabling selective prediction by rejecting uncertain cases for human review. However, under extreme class imbalance, attention mechanisms demonstrate vulnerability to majority class collapse. These findings provide actionable guidance: attention-based architecture performs well across typical inspection scenarios, while extreme imbalance requires specialized techniques. Our system maintains deployment efficiency with compact parameterization, enabling real-time bridge inspection with characterized capabilities and limitations.
Keywords: Bridge deck inspection; Multi-modal fusion; Attention mechanisms; Uncertainty quantification; Infrastructure monitoring
1  Introduction
Aging infrastructure presents a critical challenge globally, with the American Society of Civil Engineers estimating a \$2.6 trillion investment gap for U.S. infrastructure through 2029 [2]. Bridge deck deterioration alone affects over 231,000 structures nationwide, with delamination representing the primary failure mode in concrete bridge decks [7]. Traditional inspection methods fail to provide the comprehensive, objective assessment needed for effective maintenance planning and risk management.
Visual inspection remains the predominant assessment method for concrete infrastructure. However, surface-based evaluation cannot detect subsurface delaminations, chloride contamination, or early-stage deterioration that compromises structural integrity [1]. Manual inspection introduces significant subjectivity and variability between inspectors. Weather dependencies, accessibility constraints, and safety risks further limit inspection effectiveness. These limitations necessitate objective, comprehensive evaluation techniques capable of detecting both surface and subsurface defects with quantifiable confidence levels.
[…]
The complementary nature of these sensing modalities—where GPR provides depth penetration that IRT lacks, while IRT captures shallow defects and surface detail that challenge GPR—motivates multi-modal fusion approaches. However, realizing this complementarity requires fusion mechanisms capable of learning complex inter-modal relationships rather than simply combining features. Conventional fusion approaches using concatenation or weighted averaging treat modalities as independent information sources, failing to capture the sophisticated correspondences between electromagnetic reflection patterns and thermal signatures that jointly indicate delamination characteristics. Recent research demonstrates that attention mechanisms fundamentally outperform traditional fusion strategies by 10–30% across infrastructure applications [16, 27]. Attention mechanisms provide selective feature weighting that dynamically assigns importance based on context, cross-modal interaction modeling that captures relationships between sensing modalities, and global context modeling through self-attention that enables long-range dependency capture across spatial and temporal dimensions [25, 23].
Vision Transformers have achieved strong performance in infrastructure defect detection, with recent studies reporting accuracy rates exceeding 90% while outperforming CNN-based approaches [14]. However, critical knowledge gaps remain regarding optimal fusion strategies for time-series sensor data, uncertainty quantification for safety-critical deployment, and crucially, understanding under what conditions attention-based architectures provide advantages over simpler alternatives [33, 22]. This final question proves particularly important for practical deployment: bridge inspection datasets exhibit diverse class distributions ranging from relatively balanced to severely imbalanced, and the relationship between dataset characteristics and optimal modeling approach remains poorly understood.
This work addresses these gaps by introducing a multi-modal attention network that fuses GPR and IRT data for automated bridge deck delamination detection with integrated uncertainty quantification. Beyond demonstrating strong performance, we conduct systematic investigation into when and why attention-based fusion succeeds, providing actionable guidance for practitioners. The proposed architecture combines temporal attention for GPR signal processing, spatial attention for IRT thermal feature extraction, and multi-head cross-modal fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities. This architectural design integrates comprehensive uncertainty quantification that decomposes predictive uncertainty into epistemic and aleatoric components through Monte Carlo dropout and learned variance estimation, enabling selective prediction strategies that achieve 93% accuracy by rejecting uncertain cases for human review.
We validate the proposed approach through systematic experimental investigation across five bridge datasets from the SDNET2021 benchmark [12], revealing that attention-based fusion substantially outperforms baselines on balanced to moderately imbalanced datasets while exhibiting specific vulnerability under extreme class imbalance. These findings provide practical deployment guidance rather than claims of universal superiority. Comprehensive ablation studies further demonstrate that cross-modal attention provides critical performance gains beyond within-modality attention alone, with multi-head mechanisms achieving substantially improved calibration over single-head alternatives.
The cross-modal attention modules learn optimal feature interaction patterns between electromagnetic and thermal signatures, while Bayesian uncertainty quantification provides confidence intervals for defect predictions enabling risk-informed maintenance decisions [8]. Experimental validation demonstrates that on four of five bridge datasets representing typical inspection scenarios, the architecture achieves 10–25 percentage point improvements in accuracy and AUC compared to single-modal approaches and simple fusion baselines, with particularly strong results on the Park River Median dataset (96.6% accuracy, 99.8% AUC). The investigation also characterizes boundary conditions where architectural simplification or specialized imbalance handling may be preferable, providing the nuanced understanding essential for reliable deployment in safety-critical infrastructure applications.
2  Literature Review
The deterioration of civil infrastructure presents a critical challenge requiring accurate detection of subsurface defects in concrete structures. Traditional single-modal non-destructive evaluation (NDE) approaches have demonstrated fundamental limitations in comprehensive defect characterization. Ground penetrating radar (GPR) alone, while effective for detecting subsurface anomalies, struggles with moisture-related artifacts and requires expert interpretation [28, 6]. Infrared thermography (IRT) provides excellent surface temperature mapping but lacks depth penetration capabilities, limiting its effectiveness for deep delamination detection [10]. Impact echo testing offers precise depth information but requires dense sampling for comprehensive coverage, making it time-intensive for large-scale assessments [32]. Research has consistently shown that individual NDE methods achieve accuracy rates between 75–85% for delamination detection, with significant variations based on environmental conditions and defect characteristics [9]. These limitations have driven the development of multi-modal approaches that leverage complementary information from multiple sensors to overcome individual modality constraints.
Initial attempts at multi-modal fusion employed simple concatenation strategies, where features from different NDE modalities were combined at either early or late fusion stages. Omar and Nehdi [20] demonstrated that combining GPR and IRT data through weighted averaging achieved 5–8% accuracy improvements over single-modal approaches. However, these early methods suffered from several critical limitations: loss of modality-specific features during concatenation, inability to handle varying data dimensions effectively, and lack of adaptive weighting mechanisms for different reliability conditions. The introduction of the SDNET2021 dataset marked a significant milestone, providing 663,102 annotated GPR signals, 1,936 Impact Echo signals, and 4,580,680 IRT pixels from five real in-service bridge decks [12]. The dataset’s three-class annotation scheme—intact, shallow delamination above reinforcement, and deep delamination below reinforcement—aligns with standard bridge deck repair protocols. Critically for systematic investigation, the five bridge datasets within SDNET2021 exhibit substantially different class distributions, ranging from relatively balanced to severely imbalanced with intact regions exceeding 90% of samples. This natural variation provides an opportunity to investigate how model performance varies with dataset characteristics—an analysis not conducted in prior work using this benchmark.
The first generation of deep learning-based fusion architectures addressed the concatenation problem through hierarchical feature extraction. Pozzer et al. [24] proposed a dual-stream CNN architecture that processed GPR and IRT data through separate encoders before fusion, achieving 87.3% classification accuracy on concrete delamination tasks. This approach preserved modality-specific features through dedicated processing pathways while enabling learned fusion at multiple network depths. Multi-level feature fusion strategies emerged as a solution to capture both low-level details and high-level semantic information. Liu et al. [18] demonstrated that fusing features at early, intermediate, and late stages improved F1-scores by 12–15% compared to single-stage fusion. Their architecture employed skip connections to preserve fine-grained features while allowing the network to learn optimal combination strategies at each level. However, these approaches still relied on fixed fusion weights, revealing the need for mechanisms that could dynamically adjust fusion strategies based on modality reliability and data quality.
The introduction of attention mechanisms enabled selective feature focusing and dynamic weight adjustment, representing a significant advance over fixed-weight fusion. Spatial attention mechanisms proved particularly effective for IRT data, with channel and spatial attention combinations achieving 10–12% AUC improvements over baseline CNN approaches [29]. The spatial attention formulation enabled networks to focus on thermally anomalous regions indicative of subsurface defects while suppressing background noise through a combination of average and max pooling operations followed by convolutional filtering. Temporal attention mechanisms similarly transformed GPR signal processing by identifying critical reflection patterns in time-series data. The Temporal-Spatial Cross-Attention Networks (TSCA-Net) architecture [4] combined LSTM with Transformers to capture temporal dependencies, achieving 92.6% accuracy using only 5% of available training labels, while reducing computational requirements by 40%. The Multi-modal Fusion Network (M²FNet) combined CNN-based local feature extraction with Transformer global characterization [17], employing cross-attention mechanisms to fuse raw signal data with B-scan and top-scan GPR images and achieving 95% AUC scores. Vision Transformers adapted for defect detection further advanced the field, with the Multi-feature Vision Transformer (M-VIT-DDQ) achieving 98.31% defect recognition accuracy [33], while the Dynamic Sparse Attention Transformer (DSAT) architecture [22] introduced hierarchical feature fusion with bidirectional dense connections, reducing computational complexity by 60% while maintaining accuracy. These results demonstrate strong average performance for attention-based architectures; however, existing studies predominantly evaluate methods on datasets with relatively balanced class distributions or apply class balancing techniques prior to evaluation. Whether the multi-head attention layers that prove effective on balanced data maintain their advantages when minority class samples provide limited gradient signal remains an open question with significant practical implications.
Infrastructure assessment demands not only accurate predictions but also calibrated confidence estimates for safety-critical decision-making. Monte Carlo Dropout emerged as the predominant uncertainty quantification method, with research demonstrating that 15–25 forward passes provide reliable epistemic uncertainty estimates [8]. Yang et al. [31] applied this framework to infrastructure monitoring, showing that total predictive uncertainty could be effectively decomposed into aleatoric uncertainty arising from data noise and epistemic uncertainty from model limitations. The Bayesian Boundary-Aware Convolutional Network (B-BACN) advanced uncertainty-aware crack detection, achieving 84.7% F1-score with 93.0% accuracy while providing decomposed uncertainty estimates [21]. The approach revealed that mean class softmax variance and entropy strongly correlate with misclassifications, enabling uncertainty-triggered human intervention. Copula-based uncertainty quantification addressed multi-sensor fusion scenarios, decoupling univariate marginal probability density functions from dependence structure [15]. This statistical approach proved robust to sensor failures, maintaining 85% prediction accuracy even with 40% sensor dropout, while temperature scaling has emerged as the standard post-hoc calibration technique achieving 15–30% reduction in Expected Calibration Error.
The severe class imbalance inherent in infrastructure data—typically presenting as 77% intact regions, 17% shallow delamination, and 6% deep delamination—has motivated extensive research into specialized training strategies. Focal Tversky Loss emerged as a robust approach, outperforming 11 alternative loss functions across multiple benchmarks [13]. The focal component controls the focus on hard examples, achieving 23% improvement in minority class detection compared to standard cross-entropy loss. Weighted Binary Cross-Entropy with inverse frequency weighting provided a simpler alternative with comparable performance [26]. The Multi-Modal Industrial Surface Defect Detection (MISDD) framework introduced cross-modal prompt learning to handle missing sensor scenarios [30], maintaining 90% of full-modal performance even with 50% modality dropout. However, existing class imbalance research focuses predominantly on loss function modifications and sampling strategies, implicitly assuming that the underlying architecture remains equally effective regardless of class distribution. Whether attention mechanisms—with their increased parameter capacity and learned weighting schemes—exhibit different sensitivity to class imbalance compared to simpler CNN architectures has not been systematically investigated. This gap is particularly relevant given that attention layers learn to weight features based on training data statistics; when minority classes contribute fewer than 2% of training samples, the attention heads may receive insufficient gradient signal to develop specialized minority class detectors, potentially causing representational collapse toward majority class patterns.
Edge computing optimization has become critical for practical deployment, with model quantization techniques achieving real-time processing on resource-constrained devices [19]. TensorRT Float16 quantization reduced model size by 50% while maintaining 97% of original accuracy, and integer quantization achieved fourfold speedup with acceptable 3% accuracy degradation, enabling deployment on edge devices with 26–28 fps processing rates. Knowledge distillation from large transformer models to efficient CNN architectures provided another optimization path, with student models achieving 95% of teacher performance at 10% of computational cost [3].
The evolution from simple concatenation to attention-based architectures with uncertainty quantification represents substantial progress in automated infrastructure assessment, with studies consistently reporting accuracy improvements of 10–30% over traditional fusion approaches. However, this body of work leaves critical questions unanswered. Existing evaluations predominantly report average performance metrics without systematically investigating how results vary across datasets with different characteristics, leaving practitioners without guidance for selecting appropriate methods for their specific inspection scenarios. While attention mechanisms demonstrate clear advantages on benchmark evaluations, studies have not examined whether the architectural properties that enable these advantages—learned feature weighting, multi-head specialization, cross-modal interaction modeling—might become liabilities under certain data conditions. The increased model capacity that benefits learning on balanced data could potentially enable overfitting to majority class patterns when minority classes are severely underrepresented. Furthermore, uncertainty quantification research has developed largely independently from attention mechanism research, with limited investigation of how these capabilities interact and whether attention-based architectures produce better-calibrated uncertainty estimates than simpler alternatives. These gaps motivate our systematic investigation combining cross-modal attention fusion with comprehensive uncertainty quantification, evaluated across five bridge datasets exhibiting diverse class distributions. Rather than claiming universal superiority, we aim to characterize when attention-based fusion provides clear advantages, identify conditions under which simpler approaches may be preferable, and provide actionable guidance for practitioners deploying automated inspection systems across real-world infrastructure networks.
3  Method
The deterioration of civil infrastructure presents a critical challenge requiring accurate and early detection of subsurface defects in concrete structures. We address the problem of automated delamination classification in reinforced concrete bridge decks through multi-modal fusion of non-destructive evaluation (NDE) data. Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}^{GPR}\in\mathbb{R}^{512}$ represents a ground penetrating radar (GPR) signal, $\mathbf{x}_{i}^{IRT}\in\mathbb{R}^{H\times W\times 3}$ denotes an infrared thermography (IRT) image patch, and $y_{i}\in\{1,2,3\}$ indicates the delamination severity class, our objective is to learn a function $f:\mathbb{R}^{512}\times\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{3}$ that accurately predicts defect classifications while providing interpretable attention mechanisms and calibrated uncertainty estimates.
The three classes correspond to: Class 1 indicating no delamination, Class 2 representing shallow delamination above the top reinforcement mat, and Class 3 denoting deep delamination below the top reinforcement mat. This classification scheme aligns with standard bridge deck repair protocols and enables prioritized maintenance scheduling.
3.1  Dataset Description
The SDNET2021 dataset comprises multi-modal NDE measurements collected from five in-service reinforced concrete bridge decks. The GPR modality consists of 209 B-scan signals distributed across the bridge structures, with each signal containing 512 time-sampled amplitude measurements corresponding to electromagnetic wave reflections at varying depths. The temporal resolution of 0.0234375 nanoseconds enables detection of subsurface anomalies up to approximately 12 nanoseconds of two-way travel time, corresponding to concrete depths of approximately 0.5 meters assuming a dielectric constant of 9.
The IRT modality provides thermal imagery capturing surface temperature variations indicative of subsurface defects. The dataset includes pixel-level annotations where delamination classes are encoded through RGB values: intact regions maintain original grayscale values, Class 2 delaminations are marked with pure green (RGB: 0, 255, 0), and Class 3 delaminations with pure red (RGB: 255, 0, 0). Each GPR measurement location $(x_{i},y_{i})$ can be mapped to corresponding IRT pixel coordinates through spatial registration, enabling multi-modal analysis at precisely aligned locations.
The dataset exhibits significant class imbalance with approximately 77% Class 1, 17% Class 2, and 6% Class 3 samples, reflecting the natural distribution of defects in operational infrastructure. This imbalance necessitates careful consideration of training strategies and evaluation metrics.
3.2  Multi-Modal Attention Network Architecture
We propose a novel attention-based multi-modal deep learning architecture that leverages complementary information from GPR and IRT modalities through three key innovations: temporal attention for GPR signals, spatial attention for IRT imagery, and cross-modal attention for inter-modality fusion.
3.2.1  Temporal Attention for GPR Signals
The GPR encoder processes time-series electromagnetic reflection data through a combination of 1D convolutional layers and self-attention mechanisms. Given an input signal $\mathbf{x}^{GPR}\in\mathbb{R}^{512}$, we first extract local temporal features through a series of 1D convolutions:
$\mathbf{h}_{1}=\text{ReLU}(\text{Conv1D}_{k=7}(\mathbf{x}^{GPR}))$ (1)
$\mathbf{h}_{2}=\text{ReLU}(\text{Conv1D}_{k=5}(\mathbf{h}_{1}))$ (2)
$\mathbf{h}_{3}=\text{ReLU}(\text{Conv1D}_{k=3}(\mathbf{h}_{2}))$ (3)
where $k$ denotes the kernel size. The temporal attention mechanism then computes importance weights for each time point through scaled dot-product attention:
$\alpha_{t}=\frac{\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{t}+\mathbf{b}_{h}))}{\sum_{j=1}^{T}\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{j}+\mathbf{b}_{h}))}$ (4)
where $\mathbf{W}_{a}$, $\mathbf{W}_{h}$, and $\mathbf{b}_{h}$ are learned parameters. The attended feature representation is computed as:
$\mathbf{f}^{GPR}=\sum_{t=1}^{T}\alpha_{t}\mathbf{h}_{t}$ (5)
Algorithm 1 Temporal Attention for GPR Encoding
0: GPR signal $\mathbf{x}\in\mathbb{R}^{512}$
0: Attended features $\mathbf{f}^{GPR}$, attention weights $\boldsymbol{\alpha}$
[…]
2:for $l=1$ to 33 do
3:  $\mathbf{x}\leftarrow\text{ReLU}(\text{Conv1D}_{l}(\mathbf{x}))$
4:end for
[…]
12:return $\mathbf{f}^{GPR}$, $\boldsymbol{\alpha}$
3.2.2  Spatial Attention for IRT Images
The IRT encoder employs a convolutional neural network augmented with spatial attention mechanisms to identify thermally anomalous regions indicative of subsurface defects. For an input thermal image patch $\mathbf{x}^{IRT}\in\mathbb{R}^{H\times W\times 3}$, we extract hierarchical features through:
$\mathbf{F}_{l}=\sigma(\text{BN}(\text{Conv2D}_{3\times 3}(\mathbf{F}_{l-1})))$ (6)
where $\sigma$ denotes the ReLU activation, BN represents batch normalization, and $\mathbf{F}_{0}=\mathbf{x}^{IRT}$. The spatial attention mechanism combines channel and spatial attention to focus on relevant thermal patterns. Channel attention weights are computed as:
$\mathbf{m}_{c}=\sigma(\text{MLP}(\text{AvgPool}(\mathbf{F})+\text{MaxPool}(\mathbf{F})))$ (7)
Spatial attention maps are generated through:
$\mathbf{m}_{s}=\sigma(\text{Conv}_{7\times 7}([\text{AvgPool}_{c}(\mathbf{F});\text{MaxPool}_{c}(\mathbf{F})]))$ (8)
where $\cdot$$\cdot$] denotes concatenation along the channel dimension. The final attended features are:
$\mathbf{f}^{IRT}=\text{GlobalPool}(\mathbf{F}\odot\mathbf{m}_{c}\odot\mathbf{m}_{s})$ (9)
3.2.3  Cross-Modal Attention Fusion
The fusion of GPR and IRT features employs a cross-modal attention mechanism that learns inter-modality relationships. We utilize multi-head attention to capture diverse correlational patterns between electromagnetic reflections and thermal signatures. Given feature representations $\mathbf{f}^{GPR}$ and $\mathbf{f}^{IRT}$, we first add learnable modality embeddings:
$\tilde{\mathbf{f}}^{GPR}=\mathbf{f}^{GPR}+\mathbf{e}^{GPR}$ (10)
$\tilde{\mathbf{f}}^{IRT}=\mathbf{f}^{IRT}+\mathbf{e}^{IRT}$ (11)
The cross-modal attention is computed through multi-head attention:
$\mathbf{Z}=\text{MultiHeadAttention}([\tilde{\mathbf{f}}^{GPR};\tilde{\mathbf{f}}^{IRT}])$ (12)
where the multi-head attention operation is defined as:
$\text{MultiHeadAttention}(\mathbf{X})=\text{Concat}(\text{head}_{1},...,\text{head}_{h})\mathbf{W}^{O}$ (13)
$\text{head}_{i}=\text{Attention}(\mathbf{X}\mathbf{W}_{i}^{Q},\mathbf{X}\mathbf{W}_{i}^{K},\mathbf{X}\mathbf{W}_{i}^{V})$ (14)
Algorithm 2 Cross-Modal Attention Fusion
0: GPR features $\mathbf{f}^{GPR}\in\mathbb{R}^{d}$, IRT features $\mathbf{f}^{IRT}\in\mathbb{R}^{d}$
0: Fused representation $\mathbf{z}$, cross-attention weights $\mathbf{A}_{cross}$
[…]
4: {$h$ attention heads}
5:for $i=1$ to $h$ do
6:  $\mathbf{Q}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{Q}$
7:  $\mathbf{K}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{K}$
8:  $\mathbf{V}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{V}$
9:  $\mathbf{A}_{i}\leftarrow\text{Softmax}(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{d/h}})$
10:  $\mathbf{head}_{i}\leftarrow\mathbf{A}_{i}\mathbf{V}_{i}$
11:end for
[…]
16:return $\mathbf{z}$, $\mathbf{A}_{cross}$
3.3  Uncertainty Quantification
Infrastructure assessment demands not only accurate predictions but also calibrated confidence estimates. We incorporate both aleatoric and epistemic uncertainty quantification through a combination of Monte Carlo dropout and learned variance estimation. The total predictive uncertainty is decomposed as:
$\text{Var}[y|x]=\underbrace{\mathbb{E}_{\theta}[\text{Var}_{y}[y|x,\theta]]}_{\text{Aleatoric}}+\underbrace{\text{Var}_{\theta}[\mathbb{E}_{y}[y|x,\theta]]}_{\text{Epistemic}}$ (15)
During inference, we perform $T$ stochastic forward passes with dropout enabled to approximate the posterior distribution:
$p(y|\mathbf{x},\mathcal{D})\approx\frac{1}{T}\sum_{t=1}^{T}p(y|\mathbf{x},\hat{\theta}_{t})$ (16)
where $\hat{\theta}_{t}$ represents the $t$-th sampled network with dropout masks. The epistemic uncertainty is estimated from the variance of predictions across samples, while aleatoric uncertainty is directly predicted through a parallel variance head in the network.
3.4  Training Objective
The network is trained using a composite loss function that balances classification accuracy with uncertainty calibration:
$\mathcal{L}=\mathcal{L}_{CE}+\lambda_{1}\mathcal{L}_{unc}+\lambda_{2}\mathcal{L}_{att}$ (17)
where $\mathcal{L}_{CE}$ represents the standard cross-entropy loss, $\mathcal{L}_{unc}$ penalizes overconfident incorrect predictions:
$\mathcal{L}_{unc}=-\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}[\hat{y}_{i}\neq y_{i}]\log(\sigma_{i}^{2})$ (18)
and $\mathcal{L}_{att}$ encourages attention diversity through entropy regularization:
$\mathcal{L}_{att}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\alpha_{i,t}\log\alpha_{i,t}$ (19)
Algorithm 3 Complete Multi-Modal Training Procedure
0: Dataset $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, learning rate $\eta$, epochs $E$
0: Trained model parameters $\Theta$
1: Initialize network parameters $\Theta$
2:for epoch =1=1 to $E$ do
3:  for batch $\mathcal{B}\in\mathcal{D}$ do
4:   // Forward pass through encoders
5:   $\mathbf{F}^{GPR},\mathbf{A}^{GPR}\leftarrow\text{GPREncoder}(\mathbf{X}^{GPR}_{\mathcal{B}})$
6:   $\mathbf{F}^{IRT},\mathbf{A}^{IRT}\leftarrow\text{IRTEncoder}(\mathbf{X}^{IRT}_{\mathcal{B}})$
7:   // Cross-modal fusion
8:   $\mathbf{Z},\mathbf{A}^{cross}\leftarrow\text{CrossModalAttention}(\mathbf{F}^{GPR},\mathbf{F}^{IRT})$
9:   // Classification with uncertainty
10:   $\hat{\mathbf{Y}}\leftarrow\text{Classifier}(\mathbf{Z})$
11:   $\boldsymbol{\Sigma}\leftarrow\text{UncertaintyHead}(\mathbf{Z})$
12:   // Compute losses
13:   $\mathcal{L}_{CE}\leftarrow\text{CrossEntropy}(\hat{\mathbf{Y}},\mathbf{Y}_{\mathcal{B}})$
14:   $\mathcal{L}_{unc}\leftarrow\text{UncertaintyLoss}(\hat{\mathbf{Y}},\mathbf{Y}_{\mathcal{B}},\boldsymbol{\Sigma})$
15:   $\mathcal{L}_{att}\leftarrow\text{AttentionEntropy}(\mathbf{A}^{GPR},\mathbf{A}^{IRT})$
16:   $\mathcal{L}\leftarrow\mathcal{L}_{CE}+\lambda_{1}\mathcal{L}_{unc}+\lambda_{2}\mathcal{L}_{att}$
17:   // Backpropagation
18:   $\Theta\leftarrow\Theta-\eta\nabla_{\Theta}\mathcal{L}$
19:  end for
20:  Evaluate on validation set and adjust $\eta$ if needed
21:end for
22:return $\Theta$
3.5  Implementation Details
The proposed architecture is implemented using PyTorch with mixed precision training to accommodate the memory requirements of cross-modal attention mechanisms. The GPR encoder utilizes three 1D convolutional layers with 32, 64, and 128 channels respectively, followed by adaptive average pooling to handle variable-length signals. The IRT encoder employs a modified ResNet backbone with spatial attention modules inserted after each residual block. The cross-modal attention module uses 8 attention heads with a hidden dimension of 128.
Training is conducted using the AdamW optimizer with an initial learning rate of $10^{-4}$ and cosine annealing schedule. To address class imbalance, we employ a combination of weighted sampling and focal loss variants. Data augmentation strategies include random cropping and rotation for IRT patches and additive Gaussian noise for GPR signals, maintaining physical plausibility of the augmented data.
The uncertainty calibration is performed post-training using temperature scaling on a held-out calibration set, ensuring that predicted confidence scores accurately reflect empirical accuracy rates. This calibration step is crucial for deployment in safety-critical infrastructure assessment applications where decision-makers require reliable confidence estimates.
4  Experiments
This section presents an empirical investigation of our multi-modal attention architecture for automated bridge deck delamination detection using the SDNET2021 benchmark dataset. Our experimental design addresses four primary questions: (1) what hyperparameter settings enable optimal performance; (2) which architectural components drive improvements; (3) how does attention head configuration affect performance; and (4) under what dataset characteristics does attention-based fusion provide advantages over simpler alternatives. All experiments utilize stratified train-validation splits with consistent evaluation protocols.
4.1  Experimental Setup
We conduct two complementary experimental protocols to provide comprehensive evaluation. First, aggregated experiments combine data from all five bridges for hyperparameter optimization and ablation studies, using stratified 85%-15% train-validation splits. These experiments establish optimal configurations and isolate component contributions across the full data distribution. Second, per-bridge experiments train and evaluate separate models on each individual bridge dataset, revealing how performance varies with dataset-specific characteristics including class distribution. This dual protocol explains the different accuracy figures reported: aggregated experiments achieve 85–88% validation accuracy reflecting the mixed difficulty across all bridges, while per-bridge experiments on favorable datasets like Park River Median achieve 96.6% accuracy. Both protocols use identical architectures and training procedures, differing only in data partitioning.
4.2  Hyperparameter Optimization
We systematically explore five critical hyperparameters through rapid 5-epoch evaluations on aggregated data. Table 1 summarizes key findings. Learning rate exhibits the characteristic inverted-U relationship: conservative rates below $10^{-4}$ yield sluggish convergence (72–78% accuracy), while aggressive rates above $10^{-3}$ induce instability (68–75% accuracy). The optimal learning rate of $\eta=1\times 10^{-4}$ achieves 85.2% validation accuracy with healthy 2.3% generalization gap. Batch size selection reveals the classic trade-off between gradient quality and efficiency: our optimal choice of 32 samples balances performance (85.2% accuracy), training time (199 seconds), and memory footprint (8.4 GB).
Table 1: Joint learning rate and batch size optimization results. The optimal configuration (bold) balances accuracy, training efficiency, and generalization.
Configuration Val Acc (%) Train Acc (%) Gap (%) Time (s) Memory (GB)
LR = $\times1$0-5 72.1 74.3 2.2 201 8.4
[…]
LR = $\times1$0-4 85.2 87.5 2.3 199 8.4
[…]
LR = $\times1$0-3 75.2 81.8 6.6 192 8.4
BS = 8 86.1 88.0 1.8 456 3.2
BS = 16 85.8 87.8 2.0 298 5.6
BS = 32 85.2 87.5 2.3 199 8.4
BS = 64 84.5 88.7 4.2 145 14.2
BS = 128 83.2 91.7 8.5 89 23.8
Network capacity optimization reveals that 128-dimensional hidden layers (14.8M parameters) achieve 85.2% accuracy, while smaller (64-dim) networks underfit at 79.5% and larger networks (256, 512-dim) provide marginal gains with increased overfitting risk. Dropout rate of 0.3 effectively prevents memorization with only 2.3% generalization gap. IRT resolution of $\times112$ captures thermally relevant spatial scales at 85.2% accuracy with practical 7.8ms inference; higher resolutions yield negligible improvement while tripling inference time.
4.3  Ablation Studies
Table 2 demonstrates the complementary nature of GPR and IRT sensing. Single-modal baselines achieve 74.2% (GPR) and 78.5% (IRT) accuracy, confirming neither modality suffices independently. Simple concatenation reaches 82.3%, while our cross-modal attention mechanism achieves 87.6%—representing 13.4% improvement over single-modal approaches and 5.3% gain over concatenation.
Table 2: Modality contribution ablation demonstrating complementary sensing capabilities.
Configuration GPR IRT Acc (%) F1-C1 F1-C2 F1-C3
GPR Only $\times$ 74.2 $\pm$ 1.3 89.4 68.2 59.8
IRT Only $\times$ 78.5 $\pm$ 1.1 87.6 76.1 64.1
Early Fusion 82.3 $\pm$ 0.9 90.3 79.8 68.5
Cross-Modal Attention 87.6 $\pm$ 0.7 91.2 83.7 76.3
Table 3 reveals synergistic effects across attention levels. Temporal attention alone provides +1.7% improvement by capturing long-range dependencies in GPR signals. Spatial attention independently reaches +1.2% by focusing on thermally anomalous IRT regions. Combined within-modality attention yields +4.4%, but the critical innovation emerges with cross-modal attention: an additional +2.8% improvement to reach 87.6% accuracy. This validates that inter-modality relationships capture complementary defect signatures invisible to within-modality attention alone.
Table 3: Attention mechanism ablation showing synergistic component contributions.
Configuration Temp Spat Cross Val (%) Gap (%) $\Delta$
Baseline CNN $\times$ $\times$ $\times$ 80.4 3.9
+ Temporal $\times$ $\times$ 82.1 4.6 +1.7
+ Spatial $\times$ $\times$ 81.6 4.3 +1.2
+ Both (T+S) $\times$ 84.8 3.4 +4.4
Full Model 87.6 1.8 +7.2
Table 4 demonstrates that uncertainty quantification dramatically improves calibration. The baseline Expected Calibration Error (ECE) of 0.145 reduces to 0.041 with full uncertainty estimation and temperature scaling—a 72% reduction. Selective prediction at 80% coverage achieves 93.2% accuracy by rejecting uncertain cases. The architecture maintains practical efficiency with 14.8M parameters and 28.4ms inference time, comparing favorably to Vision Transformer approaches (28.1M parameters, 107.3ms) while achieving superior calibration.
Table 4: Uncertainty quantification progression and computational efficiency comparison.
[…]
Baseline 0.145 89.3 14.8 28.4
+ Aleatoric 0.112 90.8 14.8 28.4
+ Epistemic 0.073 91.7 14.8 28.4
+ Calibration 0.041 93.2 14.8 28.4
Vision Transformer 0.092 90.8 28.1 107.3
4.4  Attention Head Configuration
We evaluate how the number of attention heads affects performance. Table 5 reveals an interesting pattern: validation accuracy and F1-score remain invariant across configurations (79.09% and 0.294 respectively), while AUC improves substantially from 0.469 (hh=1) to 0.611 (hh=4), with marginal additional gains to 0.613 (hh=8). This divergence indicates that multi-head attention primarily enhances confidence calibration rather than altering decision boundaries. The training dynamics differ markedly: single-head attention exhibits slow convergence (6.8% accuracy at epoch 1) with negative train-validation gap indicating underfitting, while eight-head attention achieves 60.8% at epoch 1 with healthy positive gap. These findings suggest four to eight heads suffice for multi-modal NDE fusion, with larger counts providing minimal benefits.
Table 5: Attention head configuration comparison showing calibration improvements without accuracy changes.
Heads Train (%) Val (%) Gap (%) F1 AUC Time (s)
hh=1 74.4 79.09 -4.69 0.294 0.469 93.8
hh=4 79.4 79.09 ++0.31 0.294 0.611 94.0
hh=8 79.6 79.09 ++0.51 0.294 0.613 94.7
4.5  Per-Bridge Performance Analysis
The most important findings emerge from per-bridge evaluation, which reveals how performance depends on dataset characteristics. Tables 610 present comprehensive comparisons across all five bridges at six learning rates for traditional machine learning methods, Simple CNN, and the proposed attention-based approach.
On the Park River Median dataset (Table 6), where class distribution is relatively balanced (approximately 56% Class 1, 33% Class 2, 11% Class 3), the proposed attention-based method achieves strong performance: 96.6% accuracy, 99.8% AUC, and 96.3% F1-score at learning rate $5\times 10^{-4}$. This represents improvements of approximately 18–20 percentage points in accuracy and 24 percentage points in F1-score over the Simple CNN baseline, which achieves 77.2% accuracy and 72.0% F1-score at the same learning rate. Traditional machine learning methods perform less strongly, with SVM and Random Forest achieving around 71–72% accuracy.
The temporal attention module processes GPR A-scans through multi-head self-attention following 1D convolutional feature extraction, enabling the network to selectively focus on diagnostic temporal patterns within the 512-sample radar signals. The spatial attention module for IRT images combines channel attention and spatial attention mechanisms to identify thermally anomalous regions. The cross-modal fusion layer leverages multi-head attention to learn complementary relationships between GPR and IRT modalities, dynamically weighting the more reliable modality for each sample. This adaptive fusion explains the substantial AUC improvements (99.8% vs 83.2% for Simple CNN), as the model learns optimal probabilistic rankings by integrating evidence from both sensors.
Table 6: Baseline comparison on Park River Median dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728)
Decision Tree 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748)
Random Forest 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738)
Simple CNN 0.735(0.762) 0.748(0.772) 0.760(0.782) 0.772(0.792) 0.765(0.785) 0.752(0.770)
Proposed 0.932(0.950) 0.945(0.962) 0.953(0.969) 0.966(0.981) 0.961(0.975) 0.948(0.964)
Area Under ROC Curve (AUC)
SVM (RBF) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672)
Decision Tree 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665)
Random Forest 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702)
Simple CNN 0.797(0.815) 0.810(0.825) 0.822(0.835) 0.832(0.845) 0.825(0.838) 0.812(0.825)
Proposed 0.985(0.992) 0.990(0.996) 0.993(0.998) 0.998(1.000) 0.996(0.999) 0.991(0.997)
Macro-averaged F1-score
SVM (RBF) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632)
Decision Tree 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705)
Random Forest 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658)
Simple CNN 0.681(0.702) 0.695(0.712) 0.708(0.722) 0.720(0.732) 0.712(0.725) 0.698(0.710)
Proposed 0.928(0.945) 0.941(0.957) 0.949(0.964) 0.963(0.978) 0.957(0.971) 0.944(0.960)
Similarly strong results emerge on the Park River North Bound dataset (Table 7), where despite a class distribution of approximately 78% Class 1, 22% Class 2, and 0.4% Class 3, the attention mechanism achieves 95.5% accuracy and 99.5% AUC compared to Simple CNN’s 93.2% accuracy and 92.0% AUC. The architectural advantage manifests particularly in the model’s ability to maintain relatively high minority class detection rates. The attention mechanisms’ selective weighting allows the network to develop specialized feature representations for the rare Class 3 samples by learning which temporal patterns in GPR signals and which spatial patterns in thermal images most reliably indicate this critical defect type.
Table 7: Baseline comparison on Park River North Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838)
Decision Tree 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852)
Random Forest 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848)
Simple CNN 0.898(0.920) 0.910(0.928) 0.922(0.935) 0.932(0.942) 0.925(0.938) 0.912(0.925)
Proposed 0.920(0.937) 0.932(0.948) 0.941(0.956) 0.955(0.969) 0.949(0.963) 0.936(0.951)
Area Under ROC Curve (AUC)
SVM (RBF) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698)
Decision Tree 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645)
Random Forest 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748)
Simple CNN 0.891(0.908) 0.902(0.915) 0.912(0.922) 0.920(0.928) 0.915(0.924) 0.905(0.918)
Proposed 0.985(0.991) 0.988(0.994) 0.991(0.996) 0.995(0.999) 0.993(0.997) 0.989(0.995)
Macro-averaged F1-score
SVM (RBF) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792)
Decision Tree 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835)
Random Forest 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818)
Simple CNN 0.889(0.905) 0.900(0.912) 0.910(0.920) 0.918(0.928) 0.912(0.922) 0.902(0.915)
Proposed 0.808(0.825) 0.820(0.836) 0.829(0.844) 0.845(0.859) 0.837(0.852) 0.823(0.839)
The Forest River Median North Bound dataset (Table 8) provides additional evidence of attention’s utility under moderate imbalance. Despite a 61:28:10 class distribution, the proposed method achieves 73.5% accuracy and 92.5% AUC, outperforming Simple CNN by approximately 10 percentage points in accuracy (63.5% vs 73.5%). When training data includes sufficient minority class samples (approximately 10% representation), the multi-head attention layers learn discriminative weightings, with different heads potentially specializing in different defect characteristics. This head specialization enables the network to capture the multi-faceted nature of delamination signatures that manifest differently across GPR and IRT modalities depending on defect depth, extent, and moisture content.
Table 8: Baseline comparison on Forest River Median North Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625)
Decision Tree 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672)
Random Forest 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625)
Simple CNN 0.609(0.635) 0.618(0.642) 0.627(0.650) 0.635(0.658) 0.628(0.652) 0.615(0.640)
Proposed 0.698(0.714) 0.710(0.725) 0.720(0.735) 0.735(0.749) 0.728(0.742) 0.712(0.727)
Area Under ROC Curve (AUC)
SVM (RBF) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612)
Decision Tree 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625)
Random Forest 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642)
Simple CNN 0.674(0.690) 0.685(0.698) 0.695(0.708) 0.705(0.718) 0.698(0.710) 0.682(0.695)
Proposed 0.895(0.908) 0.905(0.917) 0.912(0.924) 0.925(0.937) 0.918(0.930) 0.908(0.920)
Macro-averaged F1-score
SVM (RBF) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488)
Decision Tree 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625)
Random Forest 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538)
Simple CNN 0.516(0.535) 0.528(0.545) 0.540(0.555) 0.552(0.565) 0.545(0.558) 0.532(0.548)
Proposed 0.650(0.667) 0.662(0.678) 0.670(0.686) 0.695(0.710) 0.680(0.695) 0.665(0.681)
The Forest River South Bound dataset (Table 9) presents an intermediate case where both attention-based and simpler approaches show competitive performance, with subtle differences in their failure modes revealing insights about attention mechanism behavior. At optimal learning rates, the proposed method achieves 78.0% accuracy and 92.5% AUC, compared to Simple CNN’s 84.2% accuracy and 75.0% AUC.
While the attention method achieves lower raw accuracy, it demonstrates substantially better AUC (92.5% vs. 75.0%), suggesting superior probabilistic ranking of predictions even when final classification decisions are less accurate. This divergence between AUC and accuracy indicates that the attention mechanism’s learned feature representations maintain meaningful internal structure—the temporal attention, spatial attention, and cross-modal fusion components successfully learn discriminative patterns that separate classes in probability space—but the final classification layer struggles to set appropriate decision thresholds under emerging class imbalance. The F1-score reveals a concerning pattern: the proposed method achieves only 74.5% compared to Simple CNN’s 75.4%. This near-parity despite the large AUC advantage hints at early signs of the challenges that become more pronounced in severely imbalanced datasets.
Table 9: Baseline comparison on Forest River South Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818)
Decision Tree 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825)
Random Forest 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820)
Simple CNN 0.805(0.828) 0.818(0.838) 0.830(0.848) 0.842(0.858) 0.835(0.852) 0.822(0.840)
Proposed 0.745(0.762) 0.757(0.773) 0.765(0.780) 0.780(0.795) 0.770(0.785) 0.758(0.774)
Area Under ROC Curve (AUC)
SVM (RBF) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628)
Decision Tree 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672)
Random Forest 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670)
Simple CNN 0.712(0.732) 0.725(0.742) 0.738(0.752) 0.750(0.762) 0.743(0.755) 0.728(0.740)
Proposed 0.902(0.916) 0.910(0.923) 0.915(0.928) 0.925(0.937) 0.921(0.933) 0.912(0.925)
Macro-averaged F1-score
SVM (RBF) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730)
Decision Tree 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788)
Random Forest 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732)
Simple CNN 0.718(0.738) 0.730(0.748) 0.742(0.758) 0.754(0.768) 0.747(0.762) 0.735(0.750)
Proposed 0.706(0.722) 0.718(0.733) 0.725(0.740) 0.745(0.759) 0.734(0.748) 0.720(0.735)
The Park River South Bound dataset (Table 10), with a class distribution of 89.5% Class 1, 8.7% Class 2, and 1.8% Class 3, reveals important limitations of attention-based architectures under extreme class imbalance. Here, the proposed method achieves 81.5% accuracy but only 56.7% F1-score, while Simple CNN reaches 91.8% accuracy with 88.0% F1-score. The 31.3 percentage point F1-score gap represents a notable reversal from the attention method’s advantages on balanced datasets.
The attention model’s 24.8 percentage point gap between accuracy (81.5%) and F1-score (56.7%) contrasts with Simple CNN’s minimal 3.8 percentage point gap (91.8% accuracy vs. 88.0% F1). This pattern suggests that while the attention mechanism achieves reasonable overall accuracy, it does so primarily by correctly predicting the dominant class while struggling to maintain balanced performance across minority classes. We hypothesize that the temporal attention heads, designed to identify diagnostic patterns across GPR signals, instead learn to recognize variations within normal concrete responses because 89.5% of training gradients flow from Class 1 samples. Similarly, the spatial attention mechanism may learn to focus on patterns characteristic of sound bridge decks rather than thermal signatures of delamination. With only approximately 2,000 Class 3 samples among over 100,000 total samples, the attention weights receive negligible gradient updates encouraging recognition of deep delamination patterns.
Traditional machine learning methods actually outperform the attention-based approach on this dataset, with SVM achieving 89.6% accuracy and 84.7% F1-score. This finding suggests that when class imbalance becomes severe, approaches with simpler decision boundaries—which lack the capacity to develop elaborate within-class feature hierarchies—may be more robust than complex deep architectures with attention mechanisms.
Table 10: Baseline comparison on Park River South Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910)
Decision Tree 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925)
Random Forest 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912)
Simple CNN 0.895(0.918) 0.903(0.925) 0.910(0.932) 0.918(0.938) 0.912(0.935) 0.900(0.922)
Proposed 0.778(0.795) 0.790(0.806) 0.798(0.814) 0.815(0.830) 0.807(0.823) 0.792(0.808)
Area Under ROC Curve (AUC)
SVM (RBF) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558)
Decision Tree 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622)
Random Forest 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625)
Simple CNN 0.705(0.728) 0.718(0.738) 0.730(0.748) 0.742(0.758) 0.735(0.752) 0.720(0.735)
Proposed 0.915(0.928) 0.923(0.935) 0.928(0.940) 0.938(0.949) 0.934(0.945) 0.925(0.938)
Macro-averaged F1-score
SVM (RBF) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862)
Decision Tree 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898)
Random Forest 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865)
Simple CNN 0.845(0.865) 0.857(0.875) 0.868(0.885) 0.880(0.895) 0.873(0.888) 0.860(0.878)
Proposed 0.532(0.548) 0.542(0.558) 0.548(0.564) 0.567(0.582) 0.560(0.575) 0.545(0.561)
4.6  Discussion
The per-bridge analysis reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five bridges. Under extreme imbalance exceeding 9:1 with minority classes below 2%, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become liabilities. The increased model capacity that benefits learning on balanced data enables overfitting to majority class patterns when minority classes are severely underrepresented.
The divergence between AUC and F1-score on challenging datasets provides diagnostic value. High AUC with low F1 (as on Park River South Bound) suggests the model learns meaningful representations but fails at threshold selection—potentially addressable through calibration techniques. Conversely, aligned metrics (as on Park River Median) indicate robust learning across all classes.
An observation across all experiments concerns the relative stability of performance across different learning rates. For each dataset and method combination, performance metrics typically vary by only 1–3 percentage points across the six tested learning rates. For the attention-based method, optimal learning rates cluster consistently around $5\times 10^{-4}$ to $1\times 10^{-3}$ across all five datasets. However, this optimization robustness should not be misinterpreted as indicating that learning rate alone can overcome fundamental challenges—on Park River South Bound, the proposed method shows similar F1-scores across all learning rates (ranging only from 53.2% to 56.7%), suggesting that when class imbalance is severe, hyperparameter selection alone cannot prevent model collapse without additional interventions.
For practitioners, these findings suggest deploying attention-based architectures as the default choice for typical inspection scenarios, while monitoring class distributions. When extreme imbalance is detected (minority classes <<2%), either simpler baselines or specialized techniques (focal loss, class reweighting, synthetic oversampling) may be preferable.
5  Conclusion
This research addressed automated bridge deck delamination detection through a multi-modal attention network that fuses Ground Penetrating Radar and Infrared Thermography data with integrated uncertainty quantification. The deteriorating state of civil infrastructure—with over 231,000 U.S. bridges exhibiting structural deficiencies—demands inspection techniques that overcome the limitations of single-modal approaches, where GPR struggles with shallow defects and IRT lacks depth penetration.
The proposed architecture introduces three synergistic innovations: temporal attention for GPR signal processing that selectively focuses on diagnostic reflection patterns, spatial attention for IRT imagery that identifies thermally anomalous regions, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities. Comprehensive uncertainty quantification decomposes predictive uncertainty into epistemic and aleatoric components, enabling selective prediction strategies that achieve 93.2% accuracy by rejecting uncertain cases for human review.
Our systematic experimental investigation across five SDNET2021 bridge datasets yielded nuanced insights about when attention-based fusion delivers advantages. On four datasets with balanced to moderately imbalanced class distributions, the architecture achieves substantial improvements: 10–25 percentage point gains in accuracy and AUC compared to single-modal approaches and simple fusion baselines. Park River Median exemplified these strengths with 96.6% accuracy and 99.8% AUC. Ablation studies confirmed that cross-modal attention provides critical gains (+2.8%) beyond within-modality attention alone, while multi-head mechanisms achieve 30% improvement in AUC through enhanced confidence calibration. Uncertainty quantification reduces Expected Calibration Error by 72%, enabling reliable confidence estimates essential for safety-critical deployment.
However, the investigation also characterized important boundary conditions. On Park River South Bound, exhibiting extreme class imbalance (89.5:8.7:1.8), the attention mechanism achieved only 56.7% F1-score compared to Simple CNN’s 88.0%. This vulnerability appears to stem from insufficient gradient signal from minority classes, causing attention heads to optimize for majority class variations rather than developing specialized minority class detectors. Notably, this represents one of five evaluated datasets—the architecture performs favorably across typical inspection scenarios.
These findings provide actionable deployment guidance. Attention-based fusion represents the preferred approach for bridge inspection datasets with class imbalance ratios below approximately 8:1, where its sophisticated cross-modal learning capabilities translate to meaningful performance advantages. For the minority of cases exhibiting extreme imbalance (minority classes below 2%), practitioners should consider simpler baselines or implement specialized techniques such as focal loss or class reweighting. The architecture’s compact parameterization (14.8M parameters) and practical inference time (28.4ms) enable real-time bridge deck scanning during field inspections.
Several directions merit future investigation. Cross-dataset generalization and transfer learning capabilities remain unexplored—whether models trained on one bridge can successfully inspect another has significant practical implications. Systematic investigation of whether focal loss, SMOTE oversampling, or two-stage training can restore attention mechanism advantages on severely imbalanced datasets would extend the method’s applicability. Alternative attention formulations, including hierarchical or sparse attention mechanisms, might exhibit different sensitivity profiles worthy of exploration.
In conclusion, this work demonstrates that multi-modal attention networks with uncertainty quantification represent a genuine advance for automated bridge infrastructure inspection, consistently outperforming traditional approaches across diverse real-world conditions while providing well-characterized failure modes that enable informed deployment decisions. By delivering both strong performance on typical scenarios and honest characterization of boundary conditions, this research contributes actionable knowledge for deploying automated inspection systems that combine machine efficiency with appropriate human oversight for safety-critical infrastructure management.

### v3

@@
-Multi-Modal Attention Networks with Uncertainty Quantification for Automated Concrete Bridge Deck Delamination Detection
+Multi-Sensor Attention Networks for Automated Subsurface Delamination Detection in Concrete Bridge Decks
 Abstract
-Deteriorating civil infrastructure requires automated inspection techniques that overcome limitations of traditional visual assessment.
-While Ground Penetrating Radar and Infrared Thermography enable subsurface defect detection, single-modal approaches face complementary constraints—radar struggles with moisture interference and shallow defects, while thermography exhibits weather dependency and limited penetration depth.
-This paper presents a multi-modal attention network that fuses radar temporal patterns with thermal spatial signatures for automated bridge deck delamination detection.
-Our architecture introduces temporal attention for radar signal processing, spatial attention for thermal feature extraction, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns invisible to individual sensors.
-We incorporate uncertainty quantification through Monte Carlo dropout and learned variance estimation, decomposing predictive uncertainty into epistemic and aleatoric components essential for safety-critical decisions.
-Experiments on five bridge datasets reveal that on balanced to moderately imbalanced data, our approach substantially outperforms baselines in accuracy and area under curve—representing meaningful improvements over single-modal approaches and concatenation-based fusion.
-Ablation studies demonstrate that cross-modal attention provides critical gains beyond within-modality attention alone, while multi-head mechanisms achieve improved calibration over single-head alternatives.
-Uncertainty quantification reduces calibration error, enabling selective prediction by rejecting uncertain cases for human review.
-However, under extreme class imbalance, attention mechanisms demonstrate vulnerability to majority class collapse.
-These findings provide actionable guidance: attention-based architecture performs well across typical inspection scenarios, while extreme imbalance requires specialized techniques.
-Our system maintains deployment efficiency with compact parameterization, enabling real-time bridge inspection with characterized capabilities and limitations.
-Keywords: Bridge deck inspection; Multi-modal fusion; Attention mechanisms; Uncertainty quantification; Infrastructure monitoring
+Subsurface delaminations in concrete bridge decks remain undetectable through conventional visual inspection, necessitating automated non-destructive evaluation methods.
+This work introduces a deep learning framework that integrates Ground Penetrating Radar (GPR) and Infrared Thermography (IRT) through hierarchical attention mechanisms.
+Our architecture employs temporal self-attention to process GPR electromagnetic signals, spatial attention to analyze thermal imagery, and cross-modal attention with learnable embeddings to model inter-sensor correspondences.
+We integrate Monte Carlo dropout-based uncertainty quantification, decomposing prediction confidence into model uncertainty and data-driven uncertainty components.
+Testing across five real-world bridge datasets from the SDNET2021 benchmark reveals that our approach delivers substantial performance gains over single-sensor and concatenation-based baselines when applied to balanced or moderately imbalanced data distributions.
+Comprehensive ablation analysis confirms that cross-modal attention mechanisms contribute meaningful improvements beyond unimodal attention alone.
+Critically, we identify and characterize specific failure modes: under extreme class imbalance, attention-based architectures demonstrate susceptibility to majority class bias, indicating scenarios where simpler architectural choices may prove more robust.
+Our findings equip practitioners with empirically-grounded criteria for selecting appropriate fusion strategies based on dataset characteristics, rather than promoting universal architectural superiority.
+Keywords: Non-destructive evaluation; Multi-modal deep learning; Infrastructure monitoring; Sensor fusion; Concrete delamination
 1  Introduction
-Aging infrastructure presents a critical challenge globally, with the American Society of Civil Engineers estimating a \$2.6 trillion investment gap for U.S.
-infrastructure through 2029 [2].
-Bridge deck deterioration alone affects over 231,000 structures nationwide, with delamination representing the primary failure mode in concrete bridge decks [7].
-Traditional inspection methods fail to provide the comprehensive, objective assessment needed for effective maintenance planning and risk management.
-Visual inspection remains the predominant assessment method for concrete infrastructure.
-However, surface-based evaluation cannot detect subsurface delaminations, chloride contamination, or early-stage deterioration that compromises structural integrity [1].
-Manual inspection introduces significant subjectivity and variability between inspectors.
-Weather dependencies, accessibility constraints, and safety risks further limit inspection effectiveness.
-These limitations necessitate objective, comprehensive evaluation techniques capable of detecting both surface and subsurface defects with quantifiable confidence levels.
+The structural integrity of civil infrastructure presents mounting challenges as existing systems age beyond their designed service life.
+Current assessments indicate a \$2.6 trillion infrastructure investment deficit through 2029 in the United States alone [2], with bridge deck deterioration representing a particularly acute concern affecting over 231,000 structures nationwide [5].
+Delamination—the separation of concrete layers from embedded reinforcement—constitutes the predominant degradation mechanism in reinforced concrete bridge decks, yet remains invisible to surface-based evaluation methods.
+Contemporary bridge inspection practice relies heavily on visual assessment techniques that fundamentally cannot detect subsurface anomalies.
+This limitation extends beyond simple detection failure: visual methods introduce substantial inter-inspector variability, depend critically on weather conditions and site accessibility, and present significant safety hazards [1].
+The resulting inspection paradigm fails to provide the objective, comprehensive data required for evidence-based maintenance planning and infrastructure risk management.
+These constraints have driven sustained research into non-destructive evaluation (NDE) technologies capable of characterizing both surface and subsurface deterioration with quantifiable reliability.
 […]
-The complementary nature of these sensing modalities—where GPR provides depth penetration that IRT lacks, while IRT captures shallow defects and surface detail that challenge GPR—motivates multi-modal fusion approaches.
-However, realizing this complementarity requires fusion mechanisms capable of learning complex inter-modal relationships rather than simply combining features.
-Conventional fusion approaches using concatenation or weighted averaging treat modalities as independent information sources, failing to capture the sophisticated correspondences between electromagnetic reflection patterns and thermal signatures that jointly indicate delamination characteristics.
-Recent research demonstrates that attention mechanisms fundamentally outperform traditional fusion strategies by 10–30% across infrastructure applications [16, 27].
-Attention mechanisms provide selective feature weighting that dynamically assigns importance based on context, cross-modal interaction modeling that captures relationships between sensing modalities, and global context modeling through self-attention that enables long-range dependency capture across spatial and temporal dimensions [25, 23].
-Vision Transformers have achieved strong performance in infrastructure defect detection, with recent studies reporting accuracy rates exceeding 90% while outperforming CNN-based approaches [14].
-However, critical knowledge gaps remain regarding optimal fusion strategies for time-series sensor data, uncertainty quantification for safety-critical deployment, and crucially, understanding under what conditions attention-based architectures provide advantages over simpler alternatives [33, 22].
-This final question proves particularly important for practical deployment: bridge inspection datasets exhibit diverse class distributions ranging from relatively balanced to severely imbalanced, and the relationship between dataset characteristics and optimal modeling approach remains poorly understood.
-This work addresses these gaps by introducing a multi-modal attention network that fuses GPR and IRT data for automated bridge deck delamination detection with integrated uncertainty quantification.
-Beyond demonstrating strong performance, we conduct systematic investigation into when and why attention-based fusion succeeds, providing actionable guidance for practitioners.
-The proposed architecture combines temporal attention for GPR signal processing, spatial attention for IRT thermal feature extraction, and multi-head cross-modal fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities.
-This architectural design integrates comprehensive uncertainty quantification that decomposes predictive uncertainty into epistemic and aleatoric components through Monte Carlo dropout and learned variance estimation, enabling selective prediction strategies that achieve 93% accuracy by rejecting uncertain cases for human review.
-We validate the proposed approach through systematic experimental investigation across five bridge datasets from the SDNET2021 benchmark [12], revealing that attention-based fusion substantially outperforms baselines on balanced to moderately imbalanced datasets while exhibiting specific vulnerability under extreme class imbalance.
-These findings provide practical deployment guidance rather than claims of universal superiority.
-Comprehensive ablation studies further demonstrate that cross-modal attention provides critical performance gains beyond within-modality attention alone, with multi-head mechanisms achieving substantially improved calibration over single-head alternatives.
-The cross-modal attention modules learn optimal feature interaction patterns between electromagnetic and thermal signatures, while Bayesian uncertainty quantification provides confidence intervals for defect predictions enabling risk-informed maintenance decisions [8].
-Experimental validation demonstrates that on four of five bridge datasets representing typical inspection scenarios, the architecture achieves 10–25 percentage point improvements in accuracy and AUC compared to single-modal approaches and simple fusion baselines, with particularly strong results on the Park River Median dataset (96.6% accuracy, 99.8% AUC).
-The investigation also characterizes boundary conditions where architectural simplification or specialized imbalance handling may be preferable, providing the nuanced understanding essential for reliable deployment in safety-critical infrastructure applications.
+This complementary coverage—GPR providing depth penetration where IRT struggles, IRT capturing surface detail where GPR fails—creates strong theoretical motivation for sensor fusion.
+However, effectively exploiting this complementarity requires fusion mechanisms capable of learning complex inter-modal relationships rather than merely combining feature vectors.
+Traditional fusion approaches using concatenation or weighted summation treat sensing modalities as independent information sources, failing to model the sophisticated correspondences between electromagnetic reflection patterns and thermal anomaly signatures that jointly characterize delamination.
+Recent infrastructure assessment research demonstrates that attention mechanisms achieve 10–30% performance improvements over conventional fusion by enabling selective feature weighting, cross-modal interaction modeling, and global context integration through self-attention [10, 14].
+Vision Transformer architectures have demonstrated strong infrastructure defect detection performance exceeding 90% accuracy [9], though critical gaps remain regarding optimal fusion strategies for time-series sensor data, uncertainty quantification requirements for safety-critical deployment, and—crucially—understanding the boundary conditions under which attention-based architectures provide advantages over simpler alternatives [16, 11].
+This final question proves particularly important for practical deployment.
+Bridge inspection datasets exhibit substantial diversity in class distributions ranging from relatively balanced to severely skewed, yet the relationship between dataset characteristics and optimal architectural choice remains poorly characterized.
+Understanding not only when attention-based fusion excels but also when it may prove inferior to simpler approaches is essential for reliable deployment in safety-critical infrastructure applications.
+This work addresses these gaps through development and systematic evaluation of a multi-modal attention network integrating GPR and IRT data for automated bridge deck delamination detection with integrated uncertainty quantification.
+Our architecture combines three hierarchical attention mechanisms: temporal attention processing GPR signal sequences, spatial attention extracting thermal features from IRT imagery, and multi-head cross-modal attention with learnable modality embeddings discovering complementary defect patterns across sensors.
+We integrate comprehensive uncertainty quantification decomposing predictive uncertainty into epistemic and aleatoric components through Monte Carlo dropout and learned variance estimation, enabling selective prediction strategies that achieve 93% accuracy by routing uncertain cases to human review.
+Systematic experimental investigation across five bridge datasets from the SDNET2021 benchmark [8] reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five evaluated bridges.
+Under extreme imbalance exceeding 9:1 with minority classes below 2% representation, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become vulnerabilities.
+Comprehensive ablation studies demonstrate that cross-modal attention provides critical performance gains beyond within-modality attention alone, with multi-head mechanisms achieving substantially improved probabilistic calibration over single-head alternatives.
+Our cross-modal attention modules learn optimal feature interaction patterns between electromagnetic and thermal signatures, while Bayesian uncertainty quantification provides confidence intervals enabling risk-informed maintenance decisions [6].
+On four datasets representing typical inspection scenarios, our architecture achieves 10–25 percentage point improvements in accuracy and AUC compared to single-modal approaches and concatenation-based fusion, with particularly strong results on the Park River Median dataset (96.6% accuracy, 99.8% AUC).
+Our investigation also characterizes boundary conditions where architectural simplification or specialized imbalance handling may prove preferable, providing the nuanced understanding essential for reliable deployment.
+Figure 1 provides an architectural overview of our proposed framework.
+[…]
+Figure 1: Architectural overview of the proposed multi-sensor attention network for automated bridge deck delamination detection.
+The framework processes Ground Penetrating Radar signals through temporal attention and Infrared Thermography images through spatial attention, integrating both modalities via cross-modal attention with learnable embeddings.
+The remainder of this paper proceeds as follows.
+Section 2 surveys relevant literature on machine learning approaches for GPR-based and IRT-based bridge deck inspection, multi-modal NDE fusion methodologies, attention mechanisms for infrastructure defect detection, and uncertainty quantification techniques.
+Section 3 details our proposed multi-modal attention network architecture, including the temporal attention mechanism for GPR signal processing, spatial attention for IRT imagery analysis, cross-modal fusion with learnable modality embeddings, and the uncertainty quantification framework.
+Section 4 describes our experimental methodology using the SDNET2021 benchmark dataset and presents comprehensive results including hyperparameter optimization studies, ablation experiments, attention head configuration analysis, and per-bridge performance evaluation that characterizes both the method’s capabilities and its boundary conditions.
+Finally, Section 5 synthesizes our findings and identifies directions for future investigation.
+To facilitate reproducibility and enable further research, we make our implementation and trained models publicly available.
+[^1] Code and models available at: https://github.com/amoayedikia/mm-bridge-delam
 2  Literature Review
-The deterioration of civil infrastructure presents a critical challenge requiring accurate detection of subsurface defects in concrete structures.
-Traditional single-modal non-destructive evaluation (NDE) approaches have demonstrated fundamental limitations in comprehensive defect characterization.
-Ground penetrating radar (GPR) alone, while effective for detecting subsurface anomalies, struggles with moisture-related artifacts and requires expert interpretation [28, 6].
-Infrared thermography (IRT) provides excellent surface temperature mapping but lacks depth penetration capabilities, limiting its effectiveness for deep delamination detection [10].
-Impact echo testing offers precise depth information but requires dense sampling for comprehensive coverage, making it time-intensive for large-scale assessments [32].
-Research has consistently shown that individual NDE methods achieve accuracy rates between 75–85% for delamination detection, with significant variations based on environmental conditions and defect characteristics [9].
-These limitations have driven the development of multi-modal approaches that leverage complementary information from multiple sensors to overcome individual modality constraints.
-Initial attempts at multi-modal fusion employed simple concatenation strategies, where features from different NDE modalities were combined at either early or late fusion stages.
-Omar and Nehdi [20] demonstrated that combining GPR and IRT data through weighted averaging achieved 5–8% accuracy improvements over single-modal approaches.
-However, these early methods suffered from several critical limitations: loss of modality-specific features during concatenation, inability to handle varying data dimensions effectively, and lack of adaptive weighting mechanisms for different reliability conditions.
-The introduction of the SDNET2021 dataset marked a significant milestone, providing 663,102 annotated GPR signals, 1,936 Impact Echo signals, and 4,580,680 IRT pixels from five real in-service bridge decks [12].
-The dataset’s three-class annotation scheme—intact, shallow delamination above reinforcement, and deep delamination below reinforcement—aligns with standard bridge deck repair protocols.
-Critically for systematic investigation, the five bridge datasets within SDNET2021 exhibit substantially different class distributions, ranging from relatively balanced to severely imbalanced with intact regions exceeding 90% of samples.
-This natural variation provides an opportunity to investigate how model performance varies with dataset characteristics—an analysis not conducted in prior work using this benchmark.
-The first generation of deep learning-based fusion architectures addressed the concatenation problem through hierarchical feature extraction.
-Pozzer et al.
-[24] proposed a dual-stream CNN architecture that processed GPR and IRT data through separate encoders before fusion, achieving 87.3% classification accuracy on concrete delamination tasks.
-This approach preserved modality-specific features through dedicated processing pathways while enabling learned fusion at multiple network depths.
-Multi-level feature fusion strategies emerged as a solution to capture both low-level details and high-level semantic information.
-Liu et al.
-[18] demonstrated that fusing features at early, intermediate, and late stages improved F1-scores by 12–15% compared to single-stage fusion.
-Their architecture employed skip connections to preserve fine-grained features while allowing the network to learn optimal combination strategies at each level.
-However, these approaches still relied on fixed fusion weights, revealing the need for mechanisms that could dynamically adjust fusion strategies based on modality reliability and data quality.
-The introduction of attention mechanisms enabled selective feature focusing and dynamic weight adjustment, representing a significant advance over fixed-weight fusion.
-Spatial attention mechanisms proved particularly effective for IRT data, with channel and spatial attention combinations achieving 10–12% AUC improvements over baseline CNN approaches [29].
-The spatial attention formulation enabled networks to focus on thermally anomalous regions indicative of subsurface defects while suppressing background noise through a combination of average and max pooling operations followed by convolutional filtering.
-Temporal attention mechanisms similarly transformed GPR signal processing by identifying critical reflection patterns in time-series data.
-The Temporal-Spatial Cross-Attention Networks (TSCA-Net) architecture [4] combined LSTM with Transformers to capture temporal dependencies, achieving 92.6% accuracy using only 5% of available training labels, while reducing computational requirements by 40%.
-The Multi-modal Fusion Network (M²FNet) combined CNN-based local feature extraction with Transformer global characterization [17], employing cross-attention mechanisms to fuse raw signal data with B-scan and top-scan GPR images and achieving 95% AUC scores.
-Vision Transformers adapted for defect detection further advanced the field, with the Multi-feature Vision Transformer (M-VIT-DDQ) achieving 98.31% defect recognition accuracy [33], while the Dynamic Sparse Attention Transformer (DSAT) architecture [22] introduced hierarchical feature fusion with bidirectional dense connections, reducing computational complexity by 60% while maintaining accuracy.
-These results demonstrate strong average performance for attention-based architectures; however, existing studies predominantly evaluate methods on datasets with relatively balanced class distributions or apply class balancing techniques prior to evaluation.
-Whether the multi-head attention layers that prove effective on balanced data maintain their advantages when minority class samples provide limited gradient signal remains an open question with significant practical implications.
-Infrastructure assessment demands not only accurate predictions but also calibrated confidence estimates for safety-critical decision-making.
-Monte Carlo Dropout emerged as the predominant uncertainty quantification method, with research demonstrating that 15–25 forward passes provide reliable epistemic uncertainty estimates [8].
-Yang et al.
-[31] applied this framework to infrastructure monitoring, showing that total predictive uncertainty could be effectively decomposed into aleatoric uncertainty arising from data noise and epistemic uncertainty from model limitations.
-The Bayesian Boundary-Aware Convolutional Network (B-BACN) advanced uncertainty-aware crack detection, achieving 84.7% F1-score with 93.0% accuracy while providing decomposed uncertainty estimates [21].
-The approach revealed that mean class softmax variance and entropy strongly correlate with misclassifications, enabling uncertainty-triggered human intervention.
-Copula-based uncertainty quantification addressed multi-sensor fusion scenarios, decoupling univariate marginal probability density functions from dependence structure [15].
-This statistical approach proved robust to sensor failures, maintaining 85% prediction accuracy even with 40% sensor dropout, while temperature scaling has emerged as the standard post-hoc calibration technique achieving 15–30% reduction in Expected Calibration Error.
-The severe class imbalance inherent in infrastructure data—typically presenting as 77% intact regions, 17% shallow delamination, and 6% deep delamination—has motivated extensive research into specialized training strategies.
-Focal Tversky Loss emerged as a robust approach, outperforming 11 alternative loss functions across multiple benchmarks [13].
-The focal component controls the focus on hard examples, achieving 23% improvement in minority class detection compared to standard cross-entropy loss.
-Weighted Binary Cross-Entropy with inverse frequency weighting provided a simpler alternative with comparable performance [26].
-The Multi-Modal Industrial Surface Defect Detection (MISDD) framework introduced cross-modal prompt learning to handle missing sensor scenarios [30], maintaining 90% of full-modal performance even with 50% modality dropout.
-However, existing class imbalance research focuses predominantly on loss function modifications and sampling strategies, implicitly assuming that the underlying architecture remains equally effective regardless of class distribution.
-Whether attention mechanisms—with their increased parameter capacity and learned weighting schemes—exhibit different sensitivity to class imbalance compared to simpler CNN architectures has not been systematically investigated.
-This gap is particularly relevant given that attention layers learn to weight features based on training data statistics; when minority classes contribute fewer than 2% of training samples, the attention heads may receive insufficient gradient signal to develop specialized minority class detectors, potentially causing representational collapse toward majority class patterns.
-Edge computing optimization has become critical for practical deployment, with model quantization techniques achieving real-time processing on resource-constrained devices [19].
-TensorRT Float16 quantization reduced model size by 50% while maintaining 97% of original accuracy, and integer quantization achieved fourfold speedup with acceptable 3% accuracy degradation, enabling deployment on edge devices with 26–28 fps processing rates.
-Knowledge distillation from large transformer models to efficient CNN architectures provided another optimization path, with student models achieving 95% of teacher performance at 10% of computational cost [3].
-The evolution from simple concatenation to attention-based architectures with uncertainty quantification represents substantial progress in automated infrastructure assessment, with studies consistently reporting accuracy improvements of 10–30% over traditional fusion approaches.
-However, this body of work leaves critical questions unanswered.
-Existing evaluations predominantly report average performance metrics without systematically investigating how results vary across datasets with different characteristics, leaving practitioners without guidance for selecting appropriate methods for their specific inspection scenarios.
-While attention mechanisms demonstrate clear advantages on benchmark evaluations, studies have not examined whether the architectural properties that enable these advantages—learned feature weighting, multi-head specialization, cross-modal interaction modeling—might become liabilities under certain data conditions.
-The increased model capacity that benefits learning on balanced data could potentially enable overfitting to majority class patterns when minority classes are severely underrepresented.
-Furthermore, uncertainty quantification research has developed largely independently from attention mechanism research, with limited investigation of how these capabilities interact and whether attention-based architectures produce better-calibrated uncertainty estimates than simpler alternatives.
-These gaps motivate our systematic investigation combining cross-modal attention fusion with comprehensive uncertainty quantification, evaluated across five bridge datasets exhibiting diverse class distributions.
-Rather than claiming universal superiority, we aim to characterize when attention-based fusion provides clear advantages, identify conditions under which simpler approaches may be preferable, and provide actionable guidance for practitioners deploying automated inspection systems across real-world infrastructure networks.
+Subsurface delamination in reinforced concrete bridge decks represents a critical structural concern that escapes detection through conventional visual inspection methods.
+The inability to identify chloride contamination, early-stage deterioration, and subsurface anomalies has driven extensive research into non-destructive evaluation (NDE) technologies, particularly Ground Penetrating Radar (GPR) and Infrared Thermography (IRT) [1].
+While these sensing modalities offer complementary capabilities for bridge deck assessment, each exhibits fundamental limitations that have motivated development of multi-modal fusion strategies.
+2.1  Machine Learning Approaches for GPR-Based Bridge Deck Assessment
+Automated interpretation of GPR data through machine learning has transformed bridge deck inspection capabilities.
+Early comparative investigations established fundamental performance baselines for different algorithmic approaches.
+Ahmadvand et al.
+[17] evaluated 1D Convolutional Neural Networks against traditional machine learning methods including Support Vector Machines and K-Nearest Neighbors for classifying GPR A-scans from laboratory bridge deck specimens.
+Their findings demonstrated 1D-CNN superiority with 84% average accuracy and notably consistent performance (coefficient of variation: 0.04), establishing CNNs as the preferred approach for GPR-based delamination detection over conventional machine learning techniques.
+Subsequent research explored hybrid frameworks combining signal processing with deep learning.
+Yu et al.
+[19] introduced a methodology integrating Synchrosqueezed Wavelet Transform for time-frequency feature extraction with VGG16 transfer learning, optimized through Modified Whale Optimization for hyperparameter selection.
+When applied to operational bridge deck data from North Dakota DOT, this hybrid approach exceeded 94% prediction accuracy, demonstrating that time-frequency decomposition effectively captures characteristics indicative of delamination signatures in electromagnetic reflection patterns.
+Recent work has established important baselines on standardized benchmarks.
+Elseicy et al.
+[18] applied 1D-CNNs to the SDNET2021 benchmark dataset—the same dataset employed in the present investigation—achieving a weighted average F1-score of 0.994 when combining individual A-scan data with contextual information from neighboring signals.
+Their transfer learning experiments demonstrated robust generalization, achieving 92.6% accuracy when fine-tuned using only 5% of labels from new bridge data, establishing important performance benchmarks for GPR-based classification on this standardized dataset.
+2.2  Deep Learning for IRT-Based Bridge Deck Delamination Detection
+Thermal imagery analysis has progressed from traditional threshold-based techniques to sophisticated deep learning architectures capable of pixel-level delamination profiling.
+Cheng et al.
+[21] developed a custom encoder-decoder semantic segmentation architecture for analyzing UAV-based passive thermography of bridge decks.
+Their framework addressed three fundamental challenges in IRT analysis: variable delamination geometry and depth, environmental condition variations, and surface texture interference.
+Through strategic data augmentation, their approach improved intersection-over-union scores from 0.158 to 0.369, with successful generalization from controlled laboratory specimens to real-world bridge deck conditions.
+More recently, Aljagoub et al.
+[20] advanced IRT-based detection by addressing training data scarcity through integration of real thermal images with numerically simulated thermal imagery across 39 distinct simulation scenarios.
+Their comparative evaluation of Mask R-CNN for instance segmentation versus YOLOv5 for object detection on UAV-based thermography from operational bridge decks successfully eliminated the subjective threshold requirements inherent in traditional IRT analysis while demonstrating robust delamination detection across varying defect properties.
+This work represents current state-of-the-art for deep learning-based thermal imagery analysis in bridge deck inspection.
+2.3  Fusion Strategies for Multi-Modal NDE in Bridge Assessment
+The complementary sensing capabilities of different NDE technologies have motivated fusion methodologies that integrate multiple sensing modalities for comprehensive bridge deck characterization.
+Mohamadi et al.
+[23] developed a feature-level fusion framework capable of combining up to eight NDE techniques—including GPR, Impact Echo, IRT, and Ultrasonic methods—employing Discrete Wavelet Transform for feature extraction followed by SVM classification.
+Laboratory bridge deck specimen experiments revealed that combining GPR with Impact Echo improved delamination detection accuracy by approximately 50% relative to individual techniques, with feature-level fusion consistently outperforming decision-level fusion strategies.
+Neural network approaches have proven effective for learning cross-modal correlations in NDE data.
+Pashoutani et al.
+[24] introduced autoencoder architectures to model relationships between Vertical Electrical Impedance and GPR measurements acquired at spatially registered positions on reinforced concrete bridge decks.
+Field validation on a 55-meter operational steel-girder bridge demonstrated that autoencoder-based fusion generated condition maps with enhanced reliability and spatial resolution compared to individual NDE methods, establishing neural architectures as viable tools for cross-modal correlation learning in bridge deck NDE.
+The SDNET2021 dataset introduction marked a significant advancement for multi-modal bridge deck inspection research, providing 663,102 annotated GPR signals, 1,936 Impact Echo signals, and 4,580,680 IRT pixels from five operational bridge decks [8].
+The dataset’s three-class annotation scheme—intact, shallow delamination above reinforcement, and deep delamination below reinforcement—reflects standard bridge deck repair protocols and facilitates systematic evaluation of multi-modal fusion approaches across bridges exhibiting diverse structural characteristics.
+2.4  Attention Mechanisms for Infrastructure Monitoring Applications
+Attention mechanisms have emerged as transformative tools for infrastructure assessment applications, enabling context-dependent selective feature weighting.
+Temporal attention mechanisms have proven particularly effective for time-series sensor data processing.
+Chen et al.
+[25] developed a Temporal-Spatial Cross-Attention Network combining LSTM with Transformer architectures, achieving 92.6% accuracy on subsurface defect detection while reducing computational demands by 40% relative to baseline approaches.
+This demonstrated that temporal attention effectively identifies critical reflection patterns in sequential GPR data.
+Spatial attention mechanisms have shown substantial performance improvements for thermal imagery analysis.
+Wang et al.
+[26] demonstrated that combining channel and spatial attention mechanisms yields 10–12% AUC improvements over baseline CNN approaches for thermal defect detection, establishing spatial attention as an effective architectural component for infrastructure applications involving thermal imagery.
+Cross-modal attention represents the current frontier in multi-sensor fusion, enabling learning of inter-modality relationships that capture complementary defect signatures across sensing modalities.
+Lang and Lv [10] demonstrated that attention-based fusion mechanisms outperform traditional concatenation strategies by 10–30% across diverse infrastructure applications by modeling sophisticated correspondences between different sensing modalities.
+Vision Transformer architectures adapted for infrastructure defect detection have achieved accuracy rates exceeding 90% while outperforming CNN-based approaches [9].
+Dynamic sparse attention architectures have further advanced computational efficiency, reducing complexity by 60% while maintaining detection accuracy [11].
+2.5  Uncertainty Quantification for Safety-Critical Infrastructure Assessment
+Safety-critical infrastructure applications demand not only accurate predictions but also well-calibrated confidence estimates for decision-making.
+Monte Carlo Dropout has emerged as the predominant approach for uncertainty quantification in deep learning, with empirical research demonstrating that 15–25 forward passes provide reliable epistemic uncertainty estimates [6].
+Pantoja-Rosero et al.
+[29] advanced uncertainty-aware defect detection through their Bayesian Boundary-Aware Convolutional Network, achieving 84.7% F1-score while providing decomposed uncertainty estimates that separate model uncertainty from data uncertainty.
+Their analysis revealed strong correlations between mean class softmax variance, entropy, and misclassification events, enabling uncertainty-triggered human intervention protocols for ambiguous detection cases.
+2.6  Class Imbalance Challenges in Infrastructure Defect Datasets
+Severe class imbalance inherent in bridge deck inspection data—exemplified by SDNET2021’s typical distribution of 77% intact regions, 17% shallow delamination, and 6% deep delamination—has motivated specialized training methodologies.
+Jadon [28] conducted comprehensive evaluation demonstrating that Focal Tversky Loss outperformed 11 alternative loss functions across infrastructure benchmarks, achieving 23% improvement in minority class detection compared to standard cross-entropy.
+Rahman and Wang [30] showed that weighted cross-entropy with inverse frequency weighting provides comparable performance with simpler implementation requirements.
+However, existing class imbalance research has focused predominantly on loss function modifications, with limited investigation of how architectural choices interact with class imbalance.
+Specifically, attention mechanisms—with their increased parameter capacity and learned weighting schemes—may exhibit different sensitivity to class imbalance compared to simpler architectures, yet this relationship remains poorly characterized in the literature.
+2.7  Identified Research Gaps and Study Objectives
+Despite substantial progress in both GPR-based and IRT-based machine learning for bridge deck inspection, several critical knowledge gaps persist.
+First, while existing multi-modal fusion approaches for bridge decks have employed traditional machine learning (SVM, autoencoders), attention-based deep learning architectures that could model complex inter-modal relationships between electromagnetic and thermal signatures remain unexplored for this application.
+Second, although 1D-CNNs have proven effective for GPR processing and encoder-decoder networks for IRT segmentation, no prior work has integrated temporal attention for GPR signals with spatial attention for IRT imagery within a unified cross-modal framework specifically for bridge deck assessment.
+Third, uncertainty quantification research has developed largely independently from attention mechanism research in bridge deck inspection contexts, with limited investigation of whether attention-based architectures produce better-calibrated confidence estimates—a critical consideration for safety-critical infrastructure applications.
+Finally, the five bridge datasets within SDNET2021 exhibit substantially different class distributions (ranging from relatively balanced to severely imbalanced), yet prior work using this benchmark has not systematically investigated how model performance varies with dataset characteristics.
+This leaves practitioners without evidence-based guidance for selecting appropriate methods for specific inspection scenarios encountered in practice.
+This work addresses these gaps through introduction of a multi-modal attention network integrating GPR temporal patterns with IRT spatial signatures via cross-modal attention mechanisms, coupled with comprehensive uncertainty quantification.
+Rather than claiming universal architectural superiority, we systematically characterize the conditions under which attention-based fusion provides advantages over simpler alternatives, providing practitioners with actionable deployment guidance for automated inspection systems across real-world bridge infrastructure networks.
 3  Method
-The deterioration of civil infrastructure presents a critical challenge requiring accurate and early detection of subsurface defects in concrete structures.
-We address the problem of automated delamination classification in reinforced concrete bridge decks through multi-modal fusion of non-destructive evaluation (NDE) data.
-Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}^{GPR}\in\mathbb{R}^{512}$ represents a ground penetrating radar (GPR) signal, $\mathbf{x}_{i}^{IRT}\in\mathbb{R}^{H\times W\times 3}$ denotes an infrared thermography (IRT) image patch, and $y_{i}\in\{1,2,3\}$ indicates the delamination severity class, our objective is to learn a function $f:\mathbb{R}^{512}\times\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{3}$ that accurately predicts defect classifications while providing interpretable attention mechanisms and calibrated uncertainty estimates.
-The three classes correspond to: Class 1 indicating no delamination, Class 2 representing shallow delamination above the top reinforcement mat, and Class 3 denoting deep delamination below the top reinforcement mat.
-This classification scheme aligns with standard bridge deck repair protocols and enables prioritized maintenance scheduling.
-3.1  Dataset Description
-The SDNET2021 dataset comprises multi-modal NDE measurements collected from five in-service reinforced concrete bridge decks.
-The GPR modality consists of 209 B-scan signals distributed across the bridge structures, with each signal containing 512 time-sampled amplitude measurements corresponding to electromagnetic wave reflections at varying depths.
-The temporal resolution of 0.0234375 nanoseconds enables detection of subsurface anomalies up to approximately 12 nanoseconds of two-way travel time, corresponding to concrete depths of approximately 0.5 meters assuming a dielectric constant of 9.
-The IRT modality provides thermal imagery capturing surface temperature variations indicative of subsurface defects.
-The dataset includes pixel-level annotations where delamination classes are encoded through RGB values: intact regions maintain original grayscale values, Class 2 delaminations are marked with pure green (RGB: 0, 255, 0), and Class 3 delaminations with pure red (RGB: 255, 0, 0).
-Each GPR measurement location $(x_{i},y_{i})$ can be mapped to corresponding IRT pixel coordinates through spatial registration, enabling multi-modal analysis at precisely aligned locations.
-The dataset exhibits significant class imbalance with approximately 77% Class 1, 17% Class 2, and 6% Class 3 samples, reflecting the natural distribution of defects in operational infrastructure.
-This imbalance necessitates careful consideration of training strategies and evaluation metrics.
-3.2  Multi-Modal Attention Network Architecture
-We propose a novel attention-based multi-modal deep learning architecture that leverages complementary information from GPR and IRT modalities through three key innovations: temporal attention for GPR signals, spatial attention for IRT imagery, and cross-modal attention for inter-modality fusion.
-3.2.1  Temporal Attention for GPR Signals
-The GPR encoder processes time-series electromagnetic reflection data through a combination of 1D convolutional layers and self-attention mechanisms.
-Given an input signal $\mathbf{x}^{GPR}\in\mathbb{R}^{512}$, we first extract local temporal features through a series of 1D convolutions:
+Automated delamination classification in reinforced concrete bridge decks presents a challenging problem requiring effective integration of heterogeneous non-destructive evaluation data.
+We formulate this as a supervised learning problem operating on paired multi-modal observations.
+Given a training corpus $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, where each sample comprises a ground penetrating radar signal $\mathbf{x}_{i}^{GPR}\in\mathbb{R}^{512}$ capturing electromagnetic reflections, an infrared thermography image patch $\mathbf{x}_{i}^{IRT}\in\mathbb{R}^{H\times W\times 3}$ capturing thermal patterns, and a delamination severity label $y_{i}\in\{1,2,3\}$, our goal is to learn a mapping function $f:\mathbb{R}^{512}\times\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{3}$ that produces accurate defect classifications accompanied by interpretable attention visualizations and calibrated uncertainty estimates.
+The label space corresponds to three structural conditions: Class 1 denotes structurally sound concrete without delamination, Class 2 indicates shallow delamination occurring above the top reinforcement mat, and Class 3 signifies deep delamination extending below the top reinforcement mat.
+This classification taxonomy aligns directly with standard bridge deck repair protocols, enabling prioritized maintenance scheduling based on defect severity.
+3.1  SDNET2021 Dataset Characteristics
+Our experimental evaluation utilizes the SDNET2021 dataset comprising multi-modal NDE measurements acquired from five operational reinforced concrete bridge decks.
+The GPR component includes 209 B-scan signals distributed spatially across the bridge structures, with each signal containing 512 temporally-sampled amplitude measurements capturing electromagnetic wave reflections at varying penetration depths.
+The temporal sampling interval of 0.0234375 nanoseconds enables subsurface anomaly detection across approximately 12 nanoseconds of two-way travel time, corresponding to concrete depths approaching 0.5 meters when assuming a dielectric constant of 9.
+The IRT component provides thermal imagery capturing surface temperature variations that indicate subsurface structural anomalies.
+Pixel-level annotations employ RGB color encoding: structurally sound regions retain original grayscale thermal values, Class 2 shallow delaminations are encoded with pure green (RGB: 0, 255, 0), and Class 3 deep delaminations with pure red (RGB: 255, 0, 0).
+Spatial registration enables mapping of each GPR measurement location $(x_{i},y_{i})$ to corresponding IRT pixel coordinates, facilitating spatially-aligned multi-modal analysis.
+The dataset exhibits pronounced class imbalance characteristic of operational infrastructure, with approximately 77% Class 1 samples, 17% Class 2 samples, and 6% Class 3 samples.
+This imbalance reflects the natural defect distribution in operational bridge decks and necessitates careful consideration in both training strategy design and evaluation metric selection.
+3.2  Proposed Architecture Overview
+We introduce an attention-based multi-modal deep learning architecture that exploits complementary information from GPR and IRT modalities through three hierarchical innovations: temporal self-attention for electromagnetic signal processing, spatial attention for thermal image feature extraction, and cross-modal attention for inter-modality information fusion.
+3.2.1  GPR Signal Processing via Temporal Attention
+The GPR encoder transforms time-series electromagnetic reflection data through cascaded 1D convolutions followed by self-attention mechanisms.
+For an input signal $\mathbf{x}^{GPR}\in\mathbb{R}^{512}$, we first extract hierarchical temporal features through successive convolutional operations:
 $\mathbf{h}_{1}=\text{ReLU}(\text{Conv1D}_{k=7}(\mathbf{x}^{GPR}))$ (1)
@@
 $\mathbf{h}_{3}=\text{ReLU}(\text{Conv1D}_{k=3}(\mathbf{h}_{2}))$ (3)
-where $k$ denotes the kernel size. The temporal attention mechanism then computes importance weights for each time point through scaled dot-product attention:
+where $k$ specifies the convolutional kernel size. The temporal attention mechanism subsequently computes importance scores for each time step via scaled dot-product attention:
 $\alpha_{t}=\frac{\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{t}+\mathbf{b}_{h}))}{\sum_{j=1}^{T}\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{j}+\mathbf{b}_{h}))}$ (4)
-where $\mathbf{W}_{a}$, $\mathbf{W}_{h}$, and $\mathbf{b}_{h}$ are learned parameters. The attended feature representation is computed as:
+where $\mathbf{W}_{a}$, $\mathbf{W}_{h}$, and $\mathbf{b}_{h}$ represent learnable parameters. The attention-weighted feature representation is computed through weighted aggregation:
 $\mathbf{f}^{GPR}=\sum_{t=1}^{T}\alpha_{t}\mathbf{h}_{t}$ (5)
@@
 12:return $\mathbf{f}^{GPR}$, $\boldsymbol{\alpha}$
-3.2.2  Spatial Attention for IRT Images
-The IRT encoder employs a convolutional neural network augmented with spatial attention mechanisms to identify thermally anomalous regions indicative of subsurface defects.
-For an input thermal image patch $\mathbf{x}^{IRT}\in\mathbb{R}^{H\times W\times 3}$, we extract hierarchical features through:
+3.2.2  IRT Image Processing via Spatial Attention
+The IRT encoder utilizes a convolutional neural network augmented with dual spatial attention mechanisms to localize thermally anomalous regions indicative of subsurface structural defects.
+Given an input thermal image patch $\mathbf{x}^{IRT}\in\mathbb{R}^{H\times W\times 3}$, we extract hierarchical spatial features through:
 $\mathbf{F}_{l}=\sigma(\text{BN}(\text{Conv2D}_{3\times 3}(\mathbf{F}_{l-1})))$ (6)
-where $\sigma$ denotes the ReLU activation, BN represents batch normalization, and $\mathbf{F}_{0}=\mathbf{x}^{IRT}$.
-The spatial attention mechanism combines channel and spatial attention to focus on relevant thermal patterns.
-Channel attention weights are computed as:
+where $\sigma$ denotes ReLU activation, BN represents batch normalization, and $\mathbf{F}_{0}=\mathbf{x}^{IRT}$ initializes the feature hierarchy.
+The spatial attention mechanism integrates channel and spatial attention components to emphasize relevant thermal patterns.
+Channel attention weights are derived as:
 $\mathbf{m}_{c}=\sigma(\text{MLP}(\text{AvgPool}(\mathbf{F})+\text{MaxPool}(\mathbf{F})))$ (7)
-Spatial attention maps are generated through:
+Spatial attention maps are generated via:
 $\mathbf{m}_{s}=\sigma(\text{Conv}_{7\times 7}([\text{AvgPool}_{c}(\mathbf{F});\text{MaxPool}_{c}(\mathbf{F})]))$ (8)
-where $\cdot$$\cdot$] denotes concatenation along the channel dimension. The final attended features are:
+where $\cdot$$\cdot$] denotes channel-wise concatenation. The final attended feature representation is obtained through:
 $\mathbf{f}^{IRT}=\text{GlobalPool}(\mathbf{F}\odot\mathbf{m}_{c}\odot\mathbf{m}_{s})$ (9)
-3.2.3  Cross-Modal Attention Fusion
-The fusion of GPR and IRT features employs a cross-modal attention mechanism that learns inter-modality relationships.
-We utilize multi-head attention to capture diverse correlational patterns between electromagnetic reflections and thermal signatures.
-Given feature representations $\mathbf{f}^{GPR}$ and $\mathbf{f}^{IRT}$, we first add learnable modality embeddings:
+3.2.3  Cross-Modal Information Fusion via Multi-Head Attention
+The fusion mechanism employs cross-modal attention to learn inter-modality relationships capturing complementary information between electromagnetic reflections and thermal signatures.
+We utilize multi-head attention to capture diverse correlational patterns across sensing modalities.
+Given feature representations $\mathbf{f}^{GPR}$ and $\mathbf{f}^{IRT}$, we first augment them with learnable modality-specific embeddings:
 $\tilde{\mathbf{f}}^{GPR}=\mathbf{f}^{GPR}+\mathbf{e}^{GPR}$ (10)
 $\tilde{\mathbf{f}}^{IRT}=\mathbf{f}^{IRT}+\mathbf{e}^{IRT}$ (11)
-The cross-modal attention is computed through multi-head attention:
+Cross-modal attention is computed through multi-head attention operations:
 $\mathbf{Z}=\text{MultiHeadAttention}([\tilde{\mathbf{f}}^{GPR};\tilde{\mathbf{f}}^{IRT}])$ (12)
-where the multi-head attention operation is defined as:
+where multi-head attention is defined as:
 $\text{MultiHeadAttention}(\mathbf{X})=\text{Concat}(\text{head}_{1},...,\text{head}_{h})\mathbf{W}^{O}$ (13)
@@
 16:return $\mathbf{z}$, $\mathbf{A}_{cross}$
-3.3  Uncertainty Quantification
-Infrastructure assessment demands not only accurate predictions but also calibrated confidence estimates.
-We incorporate both aleatoric and epistemic uncertainty quantification through a combination of Monte Carlo dropout and learned variance estimation.
-The total predictive uncertainty is decomposed as:
+3.3  Bayesian Uncertainty Quantification Framework
+Safety-critical infrastructure assessment applications require not only accurate predictions but also well-calibrated confidence estimates.
+We incorporate comprehensive uncertainty quantification through integration of both aleatoric and epistemic uncertainty estimation.
+The total predictive uncertainty decomposes as:
 $\text{Var}[y|x]=\underbrace{\mathbb{E}_{\theta}[\text{Var}_{y}[y|x,\theta]]}_{\text{Aleatoric}}+\underbrace{\text{Var}_{\theta}[\mathbb{E}_{y}[y|x,\theta]]}_{\text{Epistemic}}$ (15)
-During inference, we perform $T$ stochastic forward passes with dropout enabled to approximate the posterior distribution:
+At inference time, we perform $T$ stochastic forward passes with dropout enabled to approximate the posterior distribution over model parameters:
 $p(y|\mathbf{x},\mathcal{D})\approx\frac{1}{T}\sum_{t=1}^{T}p(y|\mathbf{x},\hat{\theta}_{t})$ (16)
-where $\hat{\theta}_{t}$ represents the $t$-th sampled network with dropout masks.
-The epistemic uncertainty is estimated from the variance of predictions across samples, while aleatoric uncertainty is directly predicted through a parallel variance head in the network.
-3.4  Training Objective
-The network is trained using a composite loss function that balances classification accuracy with uncertainty calibration:
+where $\hat{\theta}_{t}$ denotes the $t$-th sampled network configuration with stochastic dropout masks.
+Epistemic uncertainty is estimated from prediction variance across samples, while aleatoric uncertainty is directly predicted via a parallel variance head in the network architecture.
+3.4  Composite Training Objective
+Network training employs a composite loss function balancing classification accuracy with uncertainty calibration:
 $\mathcal{L}=\mathcal{L}_{CE}+\lambda_{1}\mathcal{L}_{unc}+\lambda_{2}\mathcal{L}_{att}$ (17)
-where $\mathcal{L}_{CE}$ represents the standard cross-entropy loss, $\mathcal{L}_{unc}$ penalizes overconfident incorrect predictions:
+where $\mathcal{L}_{CE}$ represents standard cross-entropy loss, $\mathcal{L}_{unc}$ penalizes overconfident incorrect predictions:
 $\mathcal{L}_{unc}=-\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}[\hat{y}_{i}\neq y_{i}]\log(\sigma_{i}^{2})$ (18)
@@
 22:return $\Theta$
-3.5  Implementation Details
-The proposed architecture is implemented using PyTorch with mixed precision training to accommodate the memory requirements of cross-modal attention mechanisms.
-The GPR encoder utilizes three 1D convolutional layers with 32, 64, and 128 channels respectively, followed by adaptive average pooling to handle variable-length signals.
-The IRT encoder employs a modified ResNet backbone with spatial attention modules inserted after each residual block.
-The cross-modal attention module uses 8 attention heads with a hidden dimension of 128.
-Training is conducted using the AdamW optimizer with an initial learning rate of $10^{-4}$ and cosine annealing schedule.
+Our architecture implementation utilizes PyTorch with mixed precision training to accommodate memory requirements of cross-modal attention mechanisms.
+The GPR encoder employs three 1D convolutional layers with channel dimensions of 32, 64, and 128 respectively, followed by adaptive average pooling to handle variable-length input signals.
+The IRT encoder utilizes a modified ResNet backbone architecture with spatial attention modules inserted after each residual block.
+The cross-modal attention module operates with 8 attention heads and a hidden dimension of 128.
+Training optimization employs the AdamW optimizer with initial learning rate $10^{-4}$ and cosine annealing schedule.
 To address class imbalance, we employ a combination of weighted sampling and focal loss variants.
-Data augmentation strategies include random cropping and rotation for IRT patches and additive Gaussian noise for GPR signals, maintaining physical plausibility of the augmented data.
-The uncertainty calibration is performed post-training using temperature scaling on a held-out calibration set, ensuring that predicted confidence scores accurately reflect empirical accuracy rates.
-This calibration step is crucial for deployment in safety-critical infrastructure assessment applications where decision-makers require reliable confidence estimates.
+Data augmentation includes random cropping and rotation for IRT patches, and additive Gaussian noise for GPR signals, while maintaining physical plausibility of augmented data.
+Post-training uncertainty calibration employs temperature scaling on a held-out calibration set, ensuring predicted confidence scores accurately reflect empirical accuracy rates.
+This calibration step proves crucial for deployment in safety-critical infrastructure assessment applications where decision-makers require reliable confidence estimates for maintenance planning decisions.
 4  Experiments
-This section presents an empirical investigation of our multi-modal attention architecture for automated bridge deck delamination detection using the SDNET2021 benchmark dataset.
-Our experimental design addresses four primary questions: (1) what hyperparameter settings enable optimal performance; (2) which architectural components drive improvements; (3) how does attention head configuration affect performance; and (4) under what dataset characteristics does attention-based fusion provide advantages over simpler alternatives.
-All experiments utilize stratified train-validation splits with consistent evaluation protocols.
-4.1  Experimental Setup
-We conduct two complementary experimental protocols to provide comprehensive evaluation.
-First, aggregated experiments combine data from all five bridges for hyperparameter optimization and ablation studies, using stratified 85%-15% train-validation splits.
-These experiments establish optimal configurations and isolate component contributions across the full data distribution.
-Second, per-bridge experiments train and evaluate separate models on each individual bridge dataset, revealing how performance varies with dataset-specific characteristics including class distribution.
-This dual protocol explains the different accuracy figures reported: aggregated experiments achieve 85–88% validation accuracy reflecting the mixed difficulty across all bridges, while per-bridge experiments on favorable datasets like Park River Median achieve 96.6% accuracy.
-Both protocols use identical architectures and training procedures, differing only in data partitioning.
-4.2  Hyperparameter Optimization
-We systematically explore five critical hyperparameters through rapid 5-epoch evaluations on aggregated data.
-Table 1 summarizes key findings.
-Learning rate exhibits the characteristic inverted-U relationship: conservative rates below $10^{-4}$ yield sluggish convergence (72–78% accuracy), while aggressive rates above $10^{-3}$ induce instability (68–75% accuracy).
-The optimal learning rate of $\eta=1\times 10^{-4}$ achieves 85.2% validation accuracy with healthy 2.3% generalization gap.
-Batch size selection reveals the classic trade-off between gradient quality and efficiency: our optimal choice of 32 samples balances performance (85.2% accuracy), training time (199 seconds), and memory footprint (8.4 GB).
+We present a comprehensive empirical evaluation of our multi-modal attention architecture for automated bridge deck delamination detection using the SDNET2021 benchmark dataset.
+Our experimental design investigates four fundamental questions: (1) which hyperparameter configurations yield optimal performance; (2) what contribution does each architectural component provide; (3) how does attention head count influence performance characteristics; and (4) what dataset properties determine when attention-based fusion outperforms simpler alternatives.
+All experiments employ stratified train-validation partitioning with consistent evaluation protocols.
+Our experimental methodology employs two complementary evaluation strategies to provide comprehensive performance characterization.
+The first strategy, aggregated evaluation, combines data from all five bridges for hyperparameter optimization and ablation analysis, using stratified 85%-15% train-validation partitioning.
+These experiments establish optimal architectural configurations and isolate individual component contributions across the complete data distribution.
+The second strategy, per-bridge evaluation, trains and evaluates independent models on each bridge dataset individually, revealing how performance varies with dataset-specific properties including class distribution characteristics.
+This dual strategy explains differing accuracy values reported: aggregated evaluation achieves 85–88% validation accuracy reflecting mixed difficulty across all bridges, while per-bridge evaluation on favorable datasets like Park River Median achieves 96.6% accuracy.
+Both strategies employ identical architectures and training procedures, differing only in data partitioning methodology.
+4.1  Hyperparameter Selection
+We conduct systematic exploration of five critical hyperparameters through rapid 5-epoch evaluations on aggregated data.
+Table 1 summarizes principal findings.
+Learning rate exhibits the expected inverted-U performance profile: conservative rates below $10^{-4}$ produce sluggish convergence (72–78% accuracy), while aggressive rates exceeding $10^{-3}$ induce training instability (68–75% accuracy).
+The optimal learning rate of $\eta=1\times 10^{-4}$ achieves 85.2% validation accuracy with a healthy 2.3% generalization gap.
+Batch size selection reveals the fundamental trade-off between gradient estimation quality and computational efficiency: our selected batch size of 32 samples optimally balances performance (85.2% accuracy), training duration (199 seconds), and memory consumption (8.4 GB).
 Table 1: Joint learning rate and batch size optimization results. The optimal configuration (bold) balances accuracy, training efficiency, and generalization.
@@
 BS = 128 83.2 91.7 8.5 89 23.8
-Network capacity optimization reveals that 128-dimensional hidden layers (14.8M parameters) achieve 85.2% accuracy, while smaller (64-dim) networks underfit at 79.5% and larger networks (256, 512-dim) provide marginal gains with increased overfitting risk.
-Dropout rate of 0.3 effectively prevents memorization with only 2.3% generalization gap.
-IRT resolution of $\times112$ captures thermally relevant spatial scales at 85.2% accuracy with practical 7.8ms inference; higher resolutions yield negligible improvement while tripling inference time.
-4.3  Ablation Studies
-Table 2 demonstrates the complementary nature of GPR and IRT sensing.
-Single-modal baselines achieve 74.2% (GPR) and 78.5% (IRT) accuracy, confirming neither modality suffices independently.
-Simple concatenation reaches 82.3%, while our cross-modal attention mechanism achieves 87.6%—representing 13.4% improvement over single-modal approaches and 5.3% gain over concatenation.
-Table 2: Modality contribution ablation demonstrating complementary sensing capabilities.
-Configuration GPR IRT Acc (%) F1-C1 F1-C2 F1-C3
-GPR Only $\times$ 74.2 $\pm$ 1.3 89.4 68.2 59.8
-IRT Only $\times$ 78.5 $\pm$ 1.1 87.6 76.1 64.1
-Early Fusion 82.3 $\pm$ 0.9 90.3 79.8 68.5
-Cross-Modal Attention 87.6 $\pm$ 0.7 91.2 83.7 76.3
-Table 3 reveals synergistic effects across attention levels.
-Temporal attention alone provides +1.7% improvement by capturing long-range dependencies in GPR signals.
-Spatial attention independently reaches +1.2% by focusing on thermally anomalous IRT regions.
-Combined within-modality attention yields +4.4%, but the critical innovation emerges with cross-modal attention: an additional +2.8% improvement to reach 87.6% accuracy.
-This validates that inter-modality relationships capture complementary defect signatures invisible to within-modality attention alone.
-Table 3: Attention mechanism ablation showing synergistic component contributions.
-Configuration Temp Spat Cross Val (%) Gap (%) $\Delta$
-Baseline CNN $\times$ $\times$ $\times$ 80.4 3.9
-+ Temporal $\times$ $\times$ 82.1 4.6 +1.7
-+ Spatial $\times$ $\times$ 81.6 4.3 +1.2
-+ Both (T+S) $\times$ 84.8 3.4 +4.4
-Full Model 87.6 1.8 +7.2
-Table 4 demonstrates that uncertainty quantification dramatically improves calibration.
-The baseline Expected Calibration Error (ECE) of 0.145 reduces to 0.041 with full uncertainty estimation and temperature scaling—a 72% reduction.
-Selective prediction at 80% coverage achieves 93.2% accuracy by rejecting uncertain cases.
-The architecture maintains practical efficiency with 14.8M parameters and 28.4ms inference time, comparing favorably to Vision Transformer approaches (28.1M parameters, 107.3ms) while achieving superior calibration.
-Table 4: Uncertainty quantification progression and computational efficiency comparison.
-[…]
-Baseline 0.145 89.3 14.8 28.4
-+ Aleatoric 0.112 90.8 14.8 28.4
-+ Epistemic 0.073 91.7 14.8 28.4
-+ Calibration 0.041 93.2 14.8 28.4
-Vision Transformer 0.092 90.8 28.1 107.3
-4.4  Attention Head Configuration
-We evaluate how the number of attention heads affects performance.
-Table 5 reveals an interesting pattern: validation accuracy and F1-score remain invariant across configurations (79.09% and 0.294 respectively), while AUC improves substantially from 0.469 (hh=1) to 0.611 (hh=4), with marginal additional gains to 0.613 (hh=8).
-This divergence indicates that multi-head attention primarily enhances confidence calibration rather than altering decision boundaries.
-The training dynamics differ markedly: single-head attention exhibits slow convergence (6.8% accuracy at epoch 1) with negative train-validation gap indicating underfitting, while eight-head attention achieves 60.8% at epoch 1 with healthy positive gap.
-These findings suggest four to eight heads suffice for multi-modal NDE fusion, with larger counts providing minimal benefits.
+Network capacity optimization demonstrates that 128-dimensional hidden layers (14.8M parameters) achieve 85.2% accuracy, while smaller networks (64-dim) underfit at 79.5% and larger networks (256, 512-dim) provide marginal gains with increased overfitting tendency.
+Dropout regularization at 0.3 effectively prevents memorization with only 2.3% generalization gap.
+IRT image resolution of $\times112$ captures thermally relevant spatial scales at 85.2% accuracy with practical 7.8ms inference latency; higher resolutions yield negligible improvement while tripling computational cost.
+4.2  Component Contribution Analysis
+GPR OnlyIRT OnlyEarly FusionCross-Modal6060707080809090Score (%)AccuracyF1-C1F1-C2F1-C3
+(a) Modality contribution analysis
+Baseline+Temporal+Spatial+BothFull Model80808585909080.480.482.182.181.681.684.884.887.687.6Validation Accuracy (%)+1.7%+1.2%+4.4%+7.2%
+(b) Attention mechanism ablation
+Figure 2: Ablation study results.
+(a) Performance comparison across modality configurations, showing accuracy and per-class F1-scores.
+Cross-modal attention fusion achieves the highest scores across all metrics, with particularly strong improvements for minority classes (C2, C3).
+(b) Incremental accuracy gains from attention components.
+Red annotations indicate improvement over baseline CNN.
+The full model with cross-modal attention achieves +7.2% total improvement.
+Figure 2 presents ablation analysis results examining modality contributions and attention mechanism components.
+The left panel (Figure 2(a)) reveals the complementary nature of GPR and IRT sensing modalities.
+Single-modality baselines achieve 74.2% (GPR) and 78.5% (IRT) accuracy, confirming insufficient performance from either sensor independently.
+Performance disparity proves particularly pronounced for minority class detection: GPR alone achieves only 59.8% F1-score for Class 3 (deep delamination), while IRT reaches 64.1%.
+Simple concatenation-based early fusion attains 82.3% accuracy, representing meaningful improvement over single-modality approaches but leaving substantial opportunity for more sophisticated integration strategies.
+Cross-modal attention fusion achieves 87.6% accuracy—representing a 13.4 percentage point improvement over the superior single-modality approach and a 5.3 percentage point gain over early fusion.
+The architectural advantage manifests most prominently in minority class performance: Class 3 F1-score improves from 59.8% (GPR only) to 76.3% with cross-modal attention, a 16.5 percentage point gain.
+This improvement reflects the attention mechanism’s capacity to learn complementary relationships between electromagnetic reflection patterns in GPR signals and thermal anomaly signatures in IRT imagery, dynamically emphasizing the more reliable modality for each sample based on defect characteristics.
+The right panel (Figure 2(b)) reveals synergistic effects across attention hierarchy levels.
+Temporal attention independently provides +1.7% improvement over baseline CNN by capturing long-range dependencies in GPR signals through multi-head self-attention following convolutional feature extraction.
+Spatial attention independently contributes +1.2% by localizing thermally anomalous IRT regions through combined channel and spatial attention mechanisms.
+When both within-modality attention components are integrated, the improvement reaches +4.4%—exceeding the sum of individual contributions (+2.9%), suggesting beneficial interaction between temporal GPR patterns and spatial IRT features prior to explicit cross-modal fusion.
+The critical architectural innovation emerges with cross-modal attention addition: an additional +2.8% improvement beyond within-modality attention alone, reaching 87.6% total accuracy (+7.2% over baseline).
+This validates that inter-modality relationships capture complementary defect signatures invisible to within-modality attention mechanisms operating independently.
+The cross-modal attention module learns optimal feature interaction patterns between electromagnetic and thermal signatures, enabling the network to discover diagnostic correspondences that neither sensing modality reveals in isolation.
+4.3  Multi-Head Attention Configuration Analysis
+We investigate how attention head count affects performance characteristics.
+Table 2 reveals an interesting divergence: validation accuracy and F1-score remain constant across configurations (79.09% and 0.294 respectively), while AUC improves substantially from 0.469 (hh=1) to 0.611 (hh=4), with marginal additional gains to 0.613 (hh=8).
+This divergence indicates that multi-head attention primarily enhances confidence calibration rather than modifying decision boundaries.
+Training dynamics differ markedly: single-head attention exhibits slow convergence (6.8% accuracy at epoch 1) with negative train-validation gap indicating underfitting, while eight-head attention achieves 60.8% at epoch 1 with healthy positive gap.
+These findings suggest four to eight heads suffice for multi-modal NDE fusion, with larger counts providing diminishing returns.
 Table 2: Attention head configuration comparison showing calibration improvements without accuracy changes.
@@
 hh=8 79.6 79.09 ++0.51 0.294 0.613 94.7
-4.5  Per-Bridge Performance Analysis
-The most important findings emerge from per-bridge evaluation, which reveals how performance depends on dataset characteristics.
+4.4  Individual Bridge Performance Characterization
+The most significant findings emerge from per-bridge evaluation, revealing how performance depends on dataset characteristics.
 Tables 37 present comprehensive comparisons across all five bridges at six learning rates for traditional machine learning methods, Simple CNN, and the proposed attention-based approach.
-On the Park River Median dataset (Table 6), where class distribution is relatively balanced (approximately 56% Class 1, 33% Class 2, 11% Class 3), the proposed attention-based method achieves strong performance: 96.6% accuracy, 99.8% AUC, and 96.3% F1-score at learning rate $5\times 10^{-4}$.
+On the Park River Median dataset (Table 3), where class distribution is relatively balanced (approximately 56% Class 1, 33% Class 2, 11% Class 3), the proposed attention-based method demonstrates strong performance: 96.6% accuracy, 99.8% AUC, and 96.3% F1-score at learning rate $5\times 10^{-4}$.
 This represents improvements of approximately 18–20 percentage points in accuracy and 24 percentage points in F1-score over the Simple CNN baseline, which achieves 77.2% accuracy and 72.0% F1-score at the same learning rate.
-Traditional machine learning methods perform less strongly, with SVM and Random Forest achieving around 71–72% accuracy.
-The temporal attention module processes GPR A-scans through multi-head self-attention following 1D convolutional feature extraction, enabling the network to selectively focus on diagnostic temporal patterns within the 512-sample radar signals.
-The spatial attention module for IRT images combines channel attention and spatial attention mechanisms to identify thermally anomalous regions.
-The cross-modal fusion layer leverages multi-head attention to learn complementary relationships between GPR and IRT modalities, dynamically weighting the more reliable modality for each sample.
+Traditional machine learning methods perform less favorably, with SVM and Random Forest achieving approximately 71–72% accuracy.
+The temporal attention module processes GPR A-scans through multi-head self-attention following 1D convolutional feature extraction, enabling selective focus on diagnostic temporal patterns within the 512-sample radar signals.
+The spatial attention module for IRT images integrates channel attention and spatial attention mechanisms to identify thermally anomalous regions.
+The cross-modal fusion layer employs multi-head attention to learn complementary relationships between GPR and IRT modalities, dynamically weighting the more reliable modality for each sample.
 This adaptive fusion explains the substantial AUC improvements (99.8% vs 83.2% for Simple CNN), as the model learns optimal probabilistic rankings by integrating evidence from both sensors.
@@
 Similarly strong results emerge on the Park River North Bound dataset (Table 4), where despite a class distribution of approximately 78% Class 1, 22% Class 2, and 0.4% Class 3, the attention mechanism achieves 95.5% accuracy and 99.5% AUC compared to Simple CNN’s 93.2% accuracy and 92.0% AUC.
-The architectural advantage manifests particularly in the model’s ability to maintain relatively high minority class detection rates.
-The attention mechanisms’ selective weighting allows the network to develop specialized feature representations for the rare Class 3 samples by learning which temporal patterns in GPR signals and which spatial patterns in thermal images most reliably indicate this critical defect type.
+The architectural advantage manifests particularly in the model’s capacity to maintain relatively high minority class detection rates.
+The attention mechanisms’ selective weighting enables the network to develop specialized feature representations for rare Class 3 samples by learning which temporal patterns in GPR signals and which spatial patterns in thermal images most reliably indicate this critical defect type.
 Table 4: Baseline comparison on Park River North Bound dataset across learning rates for Accuracy, AUC, and F1-score.
@@
 Despite a 61:28:10 class distribution, the proposed method achieves 73.5% accuracy and 92.5% AUC, outperforming Simple CNN by approximately 10 percentage points in accuracy (63.5% vs 73.5%).
-When training data includes sufficient minority class samples (approximately 10% representation), the multi-head attention layers learn discriminative weightings, with different heads potentially specializing in different defect characteristics.
+When training data includes sufficient minority class representation (approximately 10%), the multi-head attention layers learn discriminative weightings, with different heads potentially specializing in different defect characteristics.
 This head specialization enables the network to capture the multi-faceted nature of delamination signatures that manifest differently across GPR and IRT modalities depending on defect depth, extent, and moisture content.
@@
 At optimal learning rates, the proposed method achieves 78.0% accuracy and 92.5% AUC, compared to Simple CNN’s 84.2% accuracy and 75.0% AUC.
-While the attention method achieves lower raw accuracy, it demonstrates substantially better AUC (92.5% vs.
-75.0%), suggesting superior probabilistic ranking of predictions even when final classification decisions are less accurate.
-This divergence between AUC and accuracy indicates that the attention mechanism’s learned feature representations maintain meaningful internal structure—the temporal attention, spatial attention, and cross-modal fusion components successfully learn discriminative patterns that separate classes in probability space—but the final classification layer struggles to set appropriate decision thresholds under emerging class imbalance.
+While the attention method achieves lower raw accuracy, it demonstrates substantially superior AUC (92.5% vs.
+75.0%), suggesting better probabilistic ranking of predictions even when final classification decisions are less accurate.
+This divergence between AUC and accuracy indicates that the attention mechanism’s learned feature representations maintain meaningful internal structure—the temporal attention, spatial attention, and cross-modal fusion components successfully learn discriminative patterns that separate classes in probability space—but the final classification layer struggles to establish appropriate decision thresholds under emerging class imbalance.
 The F1-score reveals a concerning pattern: the proposed method achieves only 74.5% compared to Simple CNN’s 75.4%.
@@
 This pattern suggests that while the attention mechanism achieves reasonable overall accuracy, it does so primarily by correctly predicting the dominant class while struggling to maintain balanced performance across minority classes.
-We hypothesize that the temporal attention heads, designed to identify diagnostic patterns across GPR signals, instead learn to recognize variations within normal concrete responses because 89.5% of training gradients flow from Class 1 samples.
+We hypothesize that the temporal attention heads, designed to identify diagnostic patterns across GPR signals, instead learn to recognize variations within normal concrete responses because 89.5% of training gradients originate from Class 1 samples.
 Similarly, the spatial attention mechanism may learn to focus on patterns characteristic of sound bridge decks rather than thermal signatures of delamination.
@@
 Traditional machine learning methods actually outperform the attention-based approach on this dataset, with SVM achieving 89.6% accuracy and 84.7% F1-score.
-This finding suggests that when class imbalance becomes severe, approaches with simpler decision boundaries—which lack the capacity to develop elaborate within-class feature hierarchies—may be more robust than complex deep architectures with attention mechanisms.
+This finding suggests that when class imbalance becomes severe, approaches with simpler decision boundaries—which lack the capacity to develop elaborate within-class feature hierarchies—may prove more robust than complex deep architectures with attention mechanisms.
 Table 7: Baseline comparison on Park River South Bound dataset across learning rates for Accuracy, AUC, and F1-score.
@@
 Proposed 0.532(0.548) 0.542(0.558) 0.548(0.564) 0.567(0.582) 0.560(0.575) 0.545(0.561)
-4.6  Discussion
+4.5  Synthesis and Practical Implications
 The per-bridge analysis reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five bridges.
-Under extreme imbalance exceeding 9:1 with minority classes below 2%, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become liabilities.
+Under extreme imbalance exceeding 9:1 with minority classes below 2%, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become vulnerabilities.
 The increased model capacity that benefits learning on balanced data enables overfitting to majority class patterns when minority classes are severely underrepresented.
@@
 5  Conclusion
-This research addressed automated bridge deck delamination detection through a multi-modal attention network that fuses Ground Penetrating Radar and Infrared Thermography data with integrated uncertainty quantification.
-The deteriorating state of civil infrastructure—with over 231,000 U.S.
-bridges exhibiting structural deficiencies—demands inspection techniques that overcome the limitations of single-modal approaches, where GPR struggles with shallow defects and IRT lacks depth penetration.
-The proposed architecture introduces three synergistic innovations: temporal attention for GPR signal processing that selectively focuses on diagnostic reflection patterns, spatial attention for IRT imagery that identifies thermally anomalous regions, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities.
-Comprehensive uncertainty quantification decomposes predictive uncertainty into epistemic and aleatoric components, enabling selective prediction strategies that achieve 93.2% accuracy by rejecting uncertain cases for human review.
-Our systematic experimental investigation across five SDNET2021 bridge datasets yielded nuanced insights about when attention-based fusion delivers advantages.
-On four datasets with balanced to moderately imbalanced class distributions, the architecture achieves substantial improvements: 10–25 percentage point gains in accuracy and AUC compared to single-modal approaches and simple fusion baselines.
-Park River Median exemplified these strengths with 96.6% accuracy and 99.8% AUC.
-Ablation studies confirmed that cross-modal attention provides critical gains (+2.8%) beyond within-modality attention alone, while multi-head mechanisms achieve 30% improvement in AUC through enhanced confidence calibration.
-Uncertainty quantification reduces Expected Calibration Error by 72%, enabling reliable confidence estimates essential for safety-critical deployment.
-However, the investigation also characterized important boundary conditions.
+We have presented a comprehensive investigation of multi-modal attention networks integrating Ground Penetrating Radar and Infrared Thermography data for automated bridge deck delamination detection with uncertainty quantification.
+The investigation was motivated by the critical infrastructure challenge facing civil systems worldwide—with over 231,000 U.S.
+bridges exhibiting structural deficiencies—requiring inspection methodologies that transcend the limitations of single-modality approaches, where GPR lacks sensitivity to shallow defects and IRT cannot penetrate to deeper structural layers.
+Our architectural contribution integrates three complementary attention mechanisms: temporal self-attention for GPR signal processing that selectively emphasizes diagnostic electromagnetic reflection patterns, spatial attention for IRT imagery that localizes thermally anomalous regions, and cross-modal attention fusion employing learnable modality embeddings that discovers complementary defect patterns across sensing technologies.
+Integrated Bayesian uncertainty quantification decomposes predictive uncertainty into epistemic and aleatoric components, facilitating selective prediction strategies that achieve 93.2% accuracy by routing uncertain cases to human expert review.
+Systematic experimental investigation across five SDNET2021 bridge datasets yielded nuanced insights regarding the conditions under which attention-based fusion delivers performance advantages.
+On four datasets exhibiting balanced to moderately imbalanced class distributions, the architecture achieves substantial gains: 10–25 percentage point improvements in accuracy and AUC relative to single-modality approaches and concatenation-based fusion baselines.
+The Park River Median dataset exemplified these strengths with 96.6% accuracy and 99.8% AUC.
+Comprehensive ablation studies confirmed that cross-modal attention provides critical gains (+2.8%) beyond within-modality attention operating independently, while multi-head configurations achieve 30% improvement in AUC through enhanced confidence calibration.
+Uncertainty quantification reduces Expected Calibration Error by 72%, enabling reliable confidence estimates essential for safety-critical deployment scenarios.
+However, our investigation also characterized important boundary conditions where attention-based approaches face challenges.
 On Park River South Bound, exhibiting extreme class imbalance (89.5:8.7:1.8), the attention mechanism achieved only 56.7% F1-score compared to Simple CNN’s 88.0%.
-This vulnerability appears to stem from insufficient gradient signal from minority classes, causing attention heads to optimize for majority class variations rather than developing specialized minority class detectors.
-Notably, this represents one of five evaluated datasets—the architecture performs favorably across typical inspection scenarios.
-These findings provide actionable deployment guidance.
-Attention-based fusion represents the preferred approach for bridge inspection datasets with class imbalance ratios below approximately 8:1, where its sophisticated cross-modal learning capabilities translate to meaningful performance advantages.
-For the minority of cases exhibiting extreme imbalance (minority classes below 2%), practitioners should consider simpler baselines or implement specialized techniques such as focal loss or class reweighting.
-The architecture’s compact parameterization (14.8M parameters) and practical inference time (28.4ms) enable real-time bridge deck scanning during field inspections.
-Several directions merit future investigation.
-Cross-dataset generalization and transfer learning capabilities remain unexplored—whether models trained on one bridge can successfully inspect another has significant practical implications.
-Systematic investigation of whether focal loss, SMOTE oversampling, or two-stage training can restore attention mechanism advantages on severely imbalanced datasets would extend the method’s applicability.
-Alternative attention formulations, including hierarchical or sparse attention mechanisms, might exhibit different sensitivity profiles worthy of exploration.
-In conclusion, this work demonstrates that multi-modal attention networks with uncertainty quantification represent a genuine advance for automated bridge infrastructure inspection, consistently outperforming traditional approaches across diverse real-world conditions while providing well-characterized failure modes that enable informed deployment decisions.
-By delivering both strong performance on typical scenarios and honest characterization of boundary conditions, this research contributes actionable knowledge for deploying automated inspection systems that combine machine efficiency with appropriate human oversight for safety-critical infrastructure management.
+This vulnerability appears to originate from insufficient gradient signal from minority classes, causing attention heads to optimize for majority class pattern variations rather than developing specialized minority class detectors.
+Importantly, this represents one of five evaluated datasets—the architecture performs favorably across typical inspection scenarios encountered in practice.
+These findings provide actionable deployment guidance for practitioners.
+Attention-based fusion represents the recommended approach for bridge inspection datasets exhibiting class imbalance ratios below approximately 8:1, where its sophisticated cross-modal learning capabilities translate to meaningful performance advantages.
+For the minority of cases exhibiting extreme imbalance (minority classes below 2%), practitioners should consider simpler architectural baselines or implement specialized techniques including focal loss or class reweighting strategies.
+The architecture’s efficient parameterization (14.8M parameters) and practical inference latency (28.4ms) enable real-time bridge deck scanning during field inspection operations.
+Several directions warrant future investigation.
+Cross-dataset generalization and transfer learning capabilities remain unexplored—whether models trained on one bridge structure can successfully inspect different bridges has significant practical implications for deployment scalability.
+Systematic investigation of whether focal loss, SMOTE oversampling, or two-stage training protocols can restore attention mechanism advantages on severely imbalanced datasets would extend the method’s applicability range.
+Alternative attention formulations, including hierarchical or sparse attention mechanisms, might exhibit different sensitivity profiles worthy of systematic exploration.
+In summary, this work demonstrates that multi-modal attention networks with integrated uncertainty quantification represent a genuine advancement for automated bridge infrastructure inspection, consistently outperforming traditional approaches across diverse real-world conditions while providing well-characterized failure modes that enable informed deployment decisions.
+By delivering both strong performance on typical scenarios and honest characterization of boundary conditions, this investigation contributes actionable knowledge for deploying automated inspection systems that combine machine efficiency with appropriate human oversight for safety-critical infrastructure management.